### v1

# Reward Inside the Model: A Lightweight Hidden-State Reward Model for LLM’s Best-of-N sampling

Abstract
High-quality reward models are crucial for unlocking the reasoning potential of large language models (LLMs), with best-of-N voting demonstrating significant performance gains. However, current reward models, which typically operate on the textual output of LLMs, are computationally expensive and parameter-heavy, limiting their real-world applications. We introduce the Efficient Linear Hidden State Reward (ELHSR) model - a novel, highly parameter-efficient approach that leverages the rich information embedded in LLM hidden states to address these issues. ELHSR systematically outperform baselines with less than 0.005% of the parameters of baselines, requiring only a few samples for training. ELHSR also achieves orders-of-magnitude efficiency improvement with significantly less time and fewer FLOPs per sample than baseline reward models. Moreover, ELHSR exhibits robust performance even when trained only on logits, extending its applicability to some closed-source LLMs. In addition, ELHSR can also be combined with traditional reward models to achieve additional performance gains.
1 Introduction
The recent “reasoning era” (Chen et al., 2025) in large language models (LLMs) has witnessed a remarkable surge in their ability to tackle complex reasoning tasks. Models such as GPT-o3-mini (OpenAI, 2024), DeepSeek-R1 (Guo et al., 2025), Grok 3 (xAI, 2025), Claude 3.7 Sonnet (Anthropic, 2025) and Gemini 2.5 (Google, 2025) have demonstrated remarkably impressive reasoning performance. Moreover, recent studies have highlighted the considerable potential of using reward model guidance to enhance the reasoning capabilities of LLMs, typically through best-of-N sampling due to its simplicity and effectiveness (Snell et al., 2025; Brown et al., 2024).
However, current methods generally rely on predicting reward scores based on the LLM’s generated text using independently trained neural networks - often with substantial parameter counts. However, these approaches incur significant overhead, as training and deploying large reward models consume substantial computational resources and require large amount of data during both the reward model training phase and the subsequent inference stage (Guo et al., 2025; Namgoong et al., 2024; Zhang et al., 2024). This exclusive reliance on text-based information constrains these models, necessitating large training sets and complex architectures to adequately capture the full spectrum of LLM errors.
However, during the reasoning process, the LLM possesses an inherent ‘inner knowledge’ reflecting its current certainty. Existing research demonstrates that the LLM’s internal states during reasoning contain high levels of information about answer confidence, often captured through linear representations (Zhang et al., 2025; Feucht et al., 2024; Xie et al., 2024). Our preliminary experiment in Section 3.1 also supports this observation by showing that hidden states can encode a linear representation for reasoning’s correctness with high accuracy.
Figure 1: An illustration comparing traditional reward model and ELHSR.
While prior work has utilized hidden states to improve LLM’s reasoning performance, they often focus on the final answer token (Burns et al., 2023; Azaria and Mitchell, 2023; Xie et al., 2024) or on specific answer tokens within each reasoning step (Zhang et al., 2025). This approach has limitations: relying on a single token’s hidden state may not capture the full context of the reasoning process, and defining clear reasoning steps and identifying their corresponding answer tokens can be subjective and challenging. In addition, they do not provide direct reward signals, and their performance gains are limited compared to large-scale reward models. Therefore, we propose the Efficient Linear Hidden State Reward (ELHSR) model (illustrated in Figure 1). ELHSR learns a linear gating and a linear reward projection per token. The final reward is computed as a weighted average of these token-level rewards, scaled by the gating values.
Our experiments demonstrate that ELHSR systematically outperforms baseline reward models while using fewer than 0.005% their parameters. Moreover, ELHSR model is data-efficient and can be trained using only a few instances. Critically, ELHSR achieves orders-of-magnitude improvements of efficiency in average time and FLOPs per sample compared to baselines.
In summary, this work makes three main contributions:
  • We introduce ELHSR, a novel token-level reward model based on LLM hidden states. ELHSR achieves competitive reasoning performance with orders-of-magnitude less parameters and computational cost.
  • ELHSR demonstrates exceptional data efficiency and scalability, performing well with limited data, demonstrating performance upgrade with training-time scaling and test-time scaling, generalizing to logit-only training for certain closed-source LLMs, and can be combined with conventional reward models to yield further performance improvements.
  • We analyze ELHSR’s mechanisms by providing concrete examples, illustrating the way that it scores the critical components of reasoning paths.
2 Problem Statement
A common approach to enhance LLM reasoning is best-of-N sampling (Snell et al., 2025). Let $\mathcal{V}$ be the token vocabulary and let $D$ be a distribution over problem-answer pairs $(x,y)\in\mathcal{V}^{n}\times\mathcal{Y}$, where $y$ is the hidden reference solution for question $x$. Conditioned on $x$, the base LLM $\pi$ draws $N$ reasoning paths independently (i.i.d. draws from $\pi(\cdot\mid x)$), producing $r_{i}=(r_{i,1},\dots,r_{i,m_{i}})\in\mathcal{V}^{m_{i}},\quad i=1,\dots,N,$ where $m_{i}$ is the random length of the $i$-th path. We write $\mathcal{V}^{*}=\bigcup_{\ell=0}^{\infty}\mathcal{V}^{\ell}$. Simultaneously, $\pi$ emits hidden states $h_{i}=(h_{i,1},\dots,h_{i,m_{i}})\;\in\;(\mathbb{R}^{L\times d})^{m_{i}},$ where $L$ is the number of transformer layers, $d$ is the per-layer hidden-state dimension, and each $h_{i,j}\in\mathbb{R}^{L\times d}$ is the stacked hidden state after emitting $r_{i,j}$. Define the correctness indicator $F:\mathcal{V}^{*}\times\mathcal{Y}\to\{0,1\},\quad\text{s.t.}\quad F(r,y)=1\iff r\text{ yields the reference answer }y.$ We write $(\mathbb{R}^{L\times d})^{*}=\bigcup_{k=0}^{\infty}(\mathbb{R}^{L\times d})^{k}$. A parametric reward model $R_{\theta}:(\mathbb{R}^{L\times d})^{*}\to\mathbb{R}$ assigns a scalar score to each $h_{i}$ (hence for each $r_{i}$). Note that neither $\pi$ nor $R_{\theta}$ sees $y$. At test time we select $i^{*}=\arg\max_{1\leq i\leq N}R_{\theta}(h_{i})$ and return $r_{i^{*}}$. Our goal is to choose $\theta$ to maximize the probability that $r_{i^{*}}$ is correct (expectation of $F(r_{i^{*}},y)$) under $D$:
$\max_{\theta}\;\mathbb{E}_{(x,y)\sim D}\;\mathbb{E}_{(r_{1},\dots,r_{N})\sim\pi(\cdot\mid x)}\bigl{[}F\bigl{(}r_{\arg\max_{i}R_{\theta}(h_{i})},y\bigr{)}\bigr{]}.$ (1)
3 Methods
3.1 Motivation
Figure 2: Cross-validation accuracy of the PCA+LDA pipeline for predicting correctness at each layer of Llama-3.1-8B. The results demonstrate that hidden states contain information about reasoning correctness with linear representation.
We first present a preliminary experiment demonstrating the potential for enhancing reasoning using hidden states. Recent work has shown that LLM hidden states contain information about the correctness probability of a given reasoning path, often exhibiting linear representations (Xie et al., 2024; Zhang et al., 2025; Feucht et al., 2024). To verify this, we designed a simple pipeline to predict the correctness of reasoning steps based on the LLM’s hidden states.
Our pipeline extracts hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ from the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) during reasoning on the MATH dataset (Hendrycks et al., 2021), averages the hidden states across all tokens within each reasoning path to obtain a single vector representation $\bar{h}^{\ell}$ for each layer $\ell$, then reduces $\bar{h}^{\ell}$ for each $\ell$ to 50 dimensions using Principal Component Analysis (PCA) (Wold et al., 1987), and classifies the reasoning as correct or incorrect using Linear Discriminant Analysis (LDA) (Balakrishnama and Ganapathiraju, 1998). The pipeline is linear to verify that the hidden states can encode a linear representation for confidence of the answer. For more details, please refer to Appendix C.2.
We performed 5-fold cross-validation on 3000 instances of the MATH dataset. For each layer $\ell$, we trained the pipeline on the training folds and evaluated the classification accuracy on the held-out test fold. Figure 2 presents the average cross-validation accuracy for each layer.
As shown in Figure 2, the accuracy is approximately 80% on each layer, indicating a strong signal for reasoning correctness within the hidden states. This result verifies the aforementioned theory and motivates the development of a novel, lightweight, linear reward model that directly leverages this information, as described in the following sections.
3.2 Our Approach - ELHSR Model
Therefore, building upon the observation that LLM internal representations exhibit linear properties, we designed Efficient Linear Hidden State Reward (ELHSR) to be a parameter-efficient, linear model.
Our approach, ELHSR, learns a token-level reward function directly from the concatenated hidden states. ELHSR is designed to be highly parameter-efficient.
For each token $t$ in a reasoning path, we concatenate and flatten the hidden states (after residual connection) from all $L$ layers: $h_{t}=[h^{1}_{t};h^{2}_{t};...;h^{L}_{t}]\in\mathbb{R}^{Ld}$, where $d$ is the dimension of the hidden state. We then apply a linear transformation to obtain a gating value $g_{t}$ and a token-level reward $r_{t}$:
$\begin{pmatrix}\tilde{g}_{t}\\ r_{t}\end{pmatrix}=W_{ELHSR}h_{t}+b_{ELHSR}$ (2)
where $W_{ELHSR}\in\mathbb{R}^{2\times Ld}$ is the weight matrix, $b_{ELHSR}\in\mathbb{R}^{2}$ is the bias vector, $\tilde{g}_{t}$ is the pre-activation gating value, and $r_{t}$ is the token-level reward. The gating value is then passed through a sigmoid function (Han and Moraga, 1995):
$g_{t}=\sigma(\tilde{g}_{t})=\frac{1}{1+e^{-\tilde{g}_{t}}}$ (3)
The final reward $R$ for the entire reasoning path is the weighted average of the token-level rewards, using the gating values as weights:
$R=\frac{\sum_{t=1}^{T}(g_{t}\cdot r_{t})}{\max\left(\sum_{t=1}^{T}g_{t},\epsilon\right)}$ (4)
where $T$ is the number of tokens, and $\epsilon$ is a small constant (e.g., $10^{-8}$) for numerical stability. This model only contains $O(L\times d)$ parameters, which is significantly fewer than those in traditional reward models.
Since different tokens contribute differently to the overall correctness, the gating values allow the model to assign higher weights to more important tokens. In addition, we conduct ablation experiment to show that the gating mechanism can improve performance in Appendix E.2.
We use the binary cross-entropy with logits loss (Rafailov et al., 2024) for training. Let $y\in\{0,1\}$ denote the correctness label of a reasoning path, where $y=1$ indicates a correct path and $y=0$ indicates an incorrect path. For a path with hidden state representation $h$, the model predicts a scalar logit score $R(h)\in\mathbb{R}$. The binary cross-entropy with logits loss is:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\cdot\log(\sigma(R(h_{i})))+(1-y_{i})\cdot\log(1-\sigma(R(h_{i})))\right]$ (5)
where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function.
Although binary cross-entropy with logits loss generally yields good results, our experiments detailed in Appendix E.1 demonstrate that DPO (Rafailov et al., 2023) and hinge loss (Rosasco et al., 2004) can also perform well, sometimes even surpassing cross-entropy. However, given that cross-entropy performs best on average, we adopt it as our primary loss function in the main text. A comparison study among several loss functions is provided in Appendix E.1
4 Experiments
4.1 Experimental Setup
We evaluate our method on the MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and AQuA_RAT (Ling et al., 2017) datasets. The models used for generation are Llama-3.2-3B-Instruct (Meta, 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Ministral-8B-Instruct (Mistral AI, 2024). The generation temperature is 1.0 with top_p to be 0.9. We use 6000 instances (each with 8 reasoning paths) for training and 500 instances for evaluation. When training ELHSR, we use early stopping (Yao et al., 2007) based on validation set performance to mitigate overfitting. For more details on the generation process, training and evaluation, please refer to Appendix C.3, C.4 and C.5.
We compare our method against four recent open-source reward models: EurusRM-7B (Yuan et al., 2025), Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024), Starling-7B (Zhu et al., 2024) and UltraRM-13B (Cui et al., 2024). These baselines are very parameter heavy and trained on extensive datasets, while achieving leading performance among open-source reward models. For links and brief descriptions of baselines, please refer to Appendix A. For parameter count and training data size of baselines compared against ELHSR, see Table 2 for details.
To provide a more comprehensive comparison, we also evaluated a fine-tuned version of EurusRM-7B, detailed in Appendix C.6, specifically fine-tuned on the relevant datasets and model-generated solutions used in our experiments. This is despite EurusRM-7B already being pre-trained on MATH and GSM8K, which gives it a potential advantage.
4.2 Accuracy
Table 1: Different methods’ best-of-N sampling performance on MATH, GSM8K, and AQuA_RAT datasets. Results represent accuracy (percentages omitted). Our method outperforms baselines. Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.
MATH Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 42.6 48.2 46.8 50.8 52.0 52.2 54.6 56.8 55.0 51.0 Skywork-Llama3.1-8B 43.8 48.4 48.8 52.2 52.4 53.4 56.8 59.0 61.6 52.9 Starling-7B 39.6 41.2 39.8 50.4 49.0 49.0 53.8 50.2 47.0 46.7 Ultra-13B 44.6 47.4 44.4 53.0 50.6 50.4 53.6 53.0 54.0 50.1 Fine-tuned Eurus-7B 42.8 44.6 45.8 52.0 52.8 50.4 54.2 57.4 57.0 50.8 ELHSR (ours) 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
GSM8K Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 53.0 BoN@1: 80.4 BoN@1: 79.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 69.8 76.2 82.4 87.0 89.0 90.4 87.0 90.8 90.4 84.8 Skywork-Llama3.1-8B 77.2 81.8 85.2 89.6 89.4 89.2 83.0 85.8 87.4 85.4 Starling-7B 62.4 65.6 71.4 83.0 87.2 85.4 78.4 81.2 80.0 77.2 Ultra-13B 72.6 78.6 82.4 88.0 88.6 86.8 82.6 84.6 82.0 82.9 Fine-tuned Eurus-7B 79.0 84.4 86.6 90.0 90.6 90.6 88.2 91.6 92.4 88.2 ELHSR (ours) 80.0 86.4 87.6 90.6 89.6 90.6 89.8 93.2 93.4 89.0
AQuA_RAT Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 42.0 BoN@1: 53.6 BoN@1: 56.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 44.6 42.8 44.6 58.6 60.6 56.8 48.4 34.6 27.2 46.5 Skywork-Llama3.1-8B 60.4 68.0 69.4 69.0 77.0 79.2 65.6 71.2 72.0 70.2 Starling-7B 51.4 55.0 56.6 64.2 70.8 69.8 56.4 48.6 42.8 57.3 Ultra-13B 58.4 61.6 64.8 67.4 72.8 71.8 58.4 51.4 48.0 61.6 Fine-tuned Eurus-7B 57.6 63.0 57.8 67.8 72.6 70.2 70.8 71.6 70.8 66.9 ELHSR (ours) 61.0 70.4 70.8 69.6 77.0 77.0 75.2 75.8 78.0 72.8
Table 1 reports the Best-of-N performance of each reward model on the MATH, GSM8K and AQuA_RAT datasets. Our ELHSR model consistently outperforms all baseline reward models across most BoN settings, demonstrating the effectiveness of the proposed linear architecture and hinge-loss training objective. A notable observation is that while many baseline reward models exhibit strong performance on specific subsets of the data, ELHSR demonstrates consistent and robust performance across the entire dataset. While the fine-tuned EurusRM-7B delivers respectable results, it still lags behind ELHSR in accuracy and incurs substantially higher computational costs during both training and inference. In contrast, ELHSR not only achieves the best performance but does so with a parameter footprint less than 0.005% that of the baselines and is trained on only 6,000 samples per dataset (see Section 4.3 for detailed comparison). This underscores the exceptional efficiency of our approach.
It is noteworthy that the results presented here were achieved without extensive hyperparameter tuning. Further performance gains could likely be realized through more refined optimization strategies. For example, exploring alternative loss functions, as demonstrated in Appendix Table 6, increasing the training dataset size based on the scalability analysis in Section 4.4, or training ELHSR on a subset of layers as explored in Section 4.5 (Table 3), could yield further improvements. This suggests that ELHSR has significant potential for further optimization and improvement.
Additionally, we further assessed ELHSR’s performance on code comprehension tasks using the CRUXEval-O dataset (Gu et al., 2024a) to broaden our evaluation scope. We use only 600 instances for training for this dataset. The results, detailed in Appendix D (Appendix Table 5), demonstrate that ELHSR outperforms the baselines on this task as well.
4.3 Efficiency Analysis
Figure 3: Comparison of average time and FLOPs per sample for ELHSR and baselines, averaged across different datasets and task-performing models with 1-sigma error bars using standard error of the mean. The y-axis is plotted in log-scale. The results show that ELHSR achieves orders-of-magnitude higher efficiency than the baselines. See Appendix Figure 6 for full details.
Reward Model Parameters Training data size
Eurus-7B $7.1\times 10^{9}$ $5.8\times 10^{5}\times 2$
Skywork-Llama3.1-8B $7.5\times 10^{9}$ $8.0\times 10^{4}\times 2$
Starling-7B $6.7\times 10^{9}$ $1.8\times 10^{5}\times 7$
Ultra-13B $1.3\times 10^{10}$ $7.5\times 10^{5}\times 2$
ELHSR (on Llama-3.2-3B) $1.8\times 10^{5}$ $6.0\times 10^{3}\times 8$
ELHSR (on Llama-3.1-8B) $2.7\times 10^{5}$ $6.0\times 10^{3}\times 8$
ELHSR (on Ministral-8B) $3.0\times 10^{5}$ $6.0\times 10^{3}\times 8$
Table 2: Comparison of Reward Model Parameters and Training Data Size (number of questions $\times$ number of answers). Our method is remarkably efficient in terms of parameters and data.
We demonstrate ELHSR’s significant parameter efficiency and data efficiency in Table 2. Moreover, we investigate the computational efficiency of ELHSR by comparing the average time and FLOPs required to assign a reward to each sample. Figure 3 presents a comparative analysis of ELHSR and baseline reward models across various datasets and task-performing models. The results reveal that ELHSR achieves orders-of-magnitude improvements in both time and FLOPs compared to the baselines. It is noteworthy that as shown in Section 4.4, even a much smaller training dataset leads to satisfactory performance. This substantial reduction in computational cost, combined with its parameter efficiency and data efficiency, makes ELHSR a highly practical solution for reward modeling in both resource-constrained environments and large-scale applications.
4.4 Scalability with Training Set Size and Reasoning Paths
Figure 4: ELHSR has positive scaling with increased number of training samples and with the number of reasoning paths for inference. The result is averaged across datasets and task-performing models. See Appendix Figure 7 for full results.
Scaling large language models (LLMs) can be achieved through various strategies, broadly categorized as training-time scaling and test-time scaling (Snell et al., 2025; Kaplan et al., 2020; Liu et al., 2025a, b). In this section, we investigate the scalability of ELHSR with respect to both training set size and the number of reasoning paths used during inference.
From Figure 4, our results demonstrate that ELHSR’s accuracy consistently improves with both the number of training samples and the size of the reasoning paths set. Specifically, under different BoN@k settings, we find that ELHSR’s accuracy increases monotonically with training set size. Likewise, for any given training set size, ELHSR’s accuracy also improves as the number of reasoning paths used in the BoN sampling process increases.
It is interesting to note that, as illustrated in Figure 4, the performance of ELHSR with training-time scaling has not appeared to saturate even when there are 6000 training samples, which is the one used in the experiments of Section 4.2. This suggests that further increasing the training set size may lead to even greater performance improvements, highlighting the scalability of ELHSR. This indicates that ELHSR is a compelling way to scale reward modeling and hence is a practical solution for real-world applications.
4.5 Maintaining Performance with Selective Layer Utilization for Higher Efficiency
Table 3: ELHSR’s high Best-of-N sampling performance is maintained even when trained only on layers 16, 24, 28, and 32 for Llama-3.1-8B, leading to a substantial reduction in parameters and improved efficiency.
Method MATH GSM8K AQuA_RAT Avg.
BoN@1: 47.2 BoN@1: 80.4 BoN@1: 53.6
@4 @16 @64 @4 @16 @64 @4 @16 @64
ELHSR on partial layers 55.8 59.0 60.0 89.4 90.8 92.0 69.6 76.8 77.4 74.5
ELHSR on all layers 55.4 59.4 62.6 90.6 89.6 90.6 69.6 77.0 77.0 74.6
While ELHSR leverages hidden states from all layers of the LLM, we investigated the potential for further parameter reduction by selectively utilizing hidden states from a subset of layers. Specifically, for the Llama-3.1-8B model, which comprises 32 layers (plus an embedding layer), we explored training ELHSR using only the hidden states from layers 16, 24, 28, and 32, following the setting in Azaria and Mitchell (2023). This configuration reduces the parameter count to approximately 4/33 of the full-layer ELHSR model. The results, presented in Table 3, demonstrate that performance remains comparable to the full-layer ELHSR, with certain settings even showing improved accuracy. Notably, on the GSM8K dataset, the BoN@16 and BoN@64 accuracies surpass those of the full-layer ELHSR. As shown in Table 1, the performance of the layer-selected ELHSR also consistently exceeds that of baselines with significantly larger parameter counts, further underscoring the efficiency and effectiveness of our approach.
4.6 Logit-Based Training for Closed-Source LLMs
Table 4: Training ELHSR solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency. For results on GSM8K and AQuA_RAT dataset, please refer to Appendix Table 9.
MATH Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
While ELHSR offers significant advantages in terms of efficiency and performance, making it well-suited for deployment by large model providers and end-users, a limitation arises when dealing with closed-source LLMs where hidden states are inaccessible. To overcome this constraint, we leverage the fact that some closed-source models still provide access to their logits, such as GPT-3.5-turbo (OpenAI, 2023a) and GPT-4 (Achiam et al., 2023), as illustrated in Finlayson et al. (2024) and OpenAI (2023b). We train ELHSR directly on these logits, resulting in a model with a parameter count that is orders of magnitude smaller than traditional reward models - and even smaller than ELHSR trained on all layers. This allows for effortless deployment on personal devices such as smartphones and personal computers. As demonstrated in Table 4 and Appendix Table 9, the resulting logit-based ELHSR retains strong performance. Furthermore, as shown in Table 1, its performance even surpasses that of many baselines. These results highlight the versatility and adaptability of ELHSR, making it a viable solution even when access to hidden states is restricted.
4.7 Combining ELHSR with External Reward Models
Figure 5: Average performance of combined ELHSR and external reward models across different datasets and task-performing models, for both rank selection and scaled averaging. See Appendix Figure 8 for full details.
ELHSR offers a highly parameter-efficient approach to reward modeling, enabling its seamless integration with existing external reward models. Given ELHSR’s minimal parameter footprint (less than 0.005% of baseline models), the additional computational overhead introduced by incorporating ELHSR is negligible compared to relying solely on the external reward model. We explore two straightforward methods for combining ELHSR with external rewards: rank selection and scaled averaging. In rank selection, we choose the sample with the lowest (i.e. best) rank as determined by either ELHSR or the external reward model. In scaled averaging, we average the rewards from both models after normalizing them to the [0, 1] range. Detailed descriptions of these methods are provided in Appendix B.
We empirically evaluated the effectiveness of these combination strategies by integrating ELHSR with various traditional reward models across different datasets and task-performing models, assessing performance using best-of-N sampling. Figure 5 presents the average results across different datasets for both ranking and scaling methods, applied to various task-performing models. Detailed results for each dataset and task-performing model can be found in Appendix Figure 8. Our findings indicate that combining ELHSR with external reward models can lead to further improvements in accuracy, particularly on the GSM8K and AQuA-RAT datasets. The MATH dataset did not exhibit significant performance gains, potentially because ELHSR already significantly outperforms the external reward models on this benchmark. This performance boost is likely due to ELHSR and external reward models capturing different types of errors, resulting in improved overall accuracy-particularly with a higher number of best-of-N reasoning paths. This combination strategy offers a pathway to achieve further performance gains without incurring significant computational overhead.
4.8 Analysis on ELHSR
To investigate the mechanism of ELHSR, we present illustrative examples of its scoring of reasoning paths in Appendix F. The examples with the highest and lowest gating $\times$ reward values are highlighted, along with their corresponding numerical scores. Notably, ELHSR assigns relatively high or low gating $\times$ reward values to specific numerical and mathematical symbols, as well as particular tokens such as “answer”, “boxed” and the special token ‘<|eot_id|>’. This behavior suggests that ELHSR possesses the ability to identify and weigh key components within the reasoning process.
5 Related Work
Reward Modeling for Language Models’ Reasoning
The rapid advancement of Large Language Models (LLMs) in complex reasoning tasks (Wei et al., 2022; Chen et al., 2025; Hao et al., 2023; Kumar et al., 2025; Xia et al., 2025; Hao et al., 2024a) has highlighted the pivotal role of reward modeling in both training and post-training inference (Cui et al., 2024; Cao et al., 2024; Lambert et al., 2025; Sun et al., 2024; Scheurer et al., 2023; Ouyang et al., 2022). However, reward modeling faces several challenges. These include high computational costs (Guo et al., 2025; Namgoong et al., 2024; Zhang et al., 2024) and potentially suboptimal performance (Wen et al., 2025; Gao et al., 2023; Coste et al., 2024). Although prior work has proposed various approaches to mitigate these issues (Chan et al., 2024; Fisch et al., 2024; Meng et al., 2024), achieving a satisfactory balance between efficiency and performance remains a challenge. This work introduces a novel perspective by leveraging internal representations to achieve a significant improvement in efficiency with satisfactory performance.
Understanding and Leveraging the Internal Representations of Language Models
The rapid development of LLMs has spurred considerable research into their internal representations (Ferrando et al., 2024; Li et al., 2023a; Gur-Arieh et al., 2025; Pan et al., 2024; Gupta et al., 2024; Nanda et al., 2023). Hidden states of LLMs have been effectively utilized in diverse fields, including safety alignment (Zou et al., 2023; Barez et al., 2025; Zhou et al., 2024; Jiang et al., 2025; Wichers et al., 2024), faithfulness and hallucination detection (Burns et al., 2023; Li et al., 2023b; Qi et al., 2024; Ferrando et al., 2025), knowledge editing (Meng et al., 2022; Yu et al., 2024; Fang et al., 2025; Gu et al., 2024b), knowledge distillation or transfer (Dasgupta and Cohn, 2025; Wu et al., 2024) and reasoning within the latent space (Wang et al., 2025; Hao et al., 2024b). By contrast, we effectively leverage the hidden states in intermediate layers of LLMs as a direct signal for reward modeling in reasoning tasks.
6 Conclusion
In this paper, we presented the Efficient Linear Hidden State Reward (ELHSR) model, a novel reward modeling approach for large language models (LLMs) that leverages the rich information in the LLM’s internal hidden states with significant efficiency and efficacy, and can be scaled up efficiently in the training and inference process. We also demonstrated that comparable performance can be achieved by training ELHSR on a subset of layers or even on logits alone, further lowering computational costs and making it more versatile. The success of employing ELHSR in combination with standard reward models further illustrates its versatility and potential.
The linear nature of ELHSR opens up exciting avenues for future research. Its simplicity may contribute to improved interpretability of LLMs by providing insights into their reasoning process. Furthermore, the ELHSR framework could potentially be extended to multi-modal LLMs, enabling the development of efficient reward models for a broader range of applications. Additionally, future work could explore the use of ELHSR to directly guide the reinforcement learning training of LLMs. We believe that ELHSR represents a significant step towards more efficient and effective reward modeling for LLMs, paving the way for future advancements in the field.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge Mrs. Jiahe Jin from Zhiyuan College, Shanghai Jiao Tong University, for her invaluable assistance in creating the initial version of our illustrations. We also thank Mr. Zhihui Xie from The University of Hong Kong for offering precious suggestions.
Appendix Contents
[…]
Appendix A Baseline Reward Models Description
To provide further context for the baseline reward models used in our experiments, we present a brief overview and links to their respective Hugging Face Model Hub pages:
  • EurusRM-7B: https://huggingface.co/openbmb/Eurus-RM-7b. EurusRM-7B is a 7B parameter reward model trained on a mixture of UltraInteract [Yuan et al., 2025], UltraFeedback [Cui et al., 2024], and UltraSafety datasets [Guo et al., 2024], with a focus on improving reasoning performance. EurusRM-7B stands out as the best 7B RM overall and achieves similar or better performance than much larger baselines. Particularly, it outperforms GPT-4 in certain tasks.
  • Skywork-Reward-Llama-3.1-8B-v0.2: https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B-v0.2. This 8B parameter reward model is built on the Llama-3 architecture [Grattafiori et al., 2024] and trained on a curated dataset of 80K high-quality preference pairs. It excels at handling preferences in complex scenarios, including challenging preference pairs, and span various domains such as mathematics, coding, and safety. As of October 2024, Skywork-Reward-Llama-3.1-8B-v0.2 ranks first among 8B models.
  • Starling-RM-7B: https://huggingface.co/berkeley-nest/Starling-RM-7B-alpha. Starling-RM-7B is a reward model trained from Llama2-7B-Chat [Touvron et al., 2023], following the method of training reward models in the instructGPT paper [Ouyang et al., 2022]. It is trained with the Nectar preference dataset [Zhu et al., 2024], which is based on GPT-4’s preferences.
  • UltraRM-13B: https://huggingface.co/openbmb/UltraRM-13b. UltraRM is a reward model initialized by Llama2-13B [Touvron et al., 2023] and fine-tuned on the UltraFeedback dataset [Cui et al., 2024]. It achieves state-of-the-art performance among open-source reward models on public preference test sets.
Appendix B Detailed Approach for Combining ELHSR with External Reward Models
In the Best-of-N (BoN) setting, we explore two strategies for combining the ELHSR reward ($R_{ELHSR}$) with an external reward signal ($R_{ext}$), typically from a traditional reward model. We have $N$ reasoning paths.
  1. 1.
    Ranking-based Combination: Let $rank_{ELHSR}(i)$ and $rank_{ext}(i)$ be the ranks of the $i$-th reasoning path according to ELHSR and the external reward model, respectively (lower rank is better). The relative rank is:
    $rank_{rel}(i)=\max(rank_{ELHSR}(i),rank_{ext}(i))$ (6)
    We select the path with the lowest relative rank:
    $i^{*}=\arg\min_{i\in\{1,...,N\}}rank_{rel}(i)$ (7)
  2. 2.
    Scaling-based Combination: We linearly scale both rewards to the range [0, 1] for the $N$ paths [^1] If all the rewards are equal, we set them all to $1.0$, but this is almost impossible in practice.:
    $\tilde{R}_{ELHSR}(i)=\frac{R_{ELHSR}(i)-\min_{j}R_{ELHSR}(j)}{\max_{j}R_{ELHSR}(j)-\min_{j}R_{ELHSR}(j)}$ (8)
    $\tilde{R}_{ext}(i)=\frac{R_{ext}(i)-\min_{j}R_{ext}(j)}{\max_{j}R_{ext}(j)-\min_{j}R_{ext}(j)}$ (9)
    The aggregated reward is:
    $R_{agg}(i)=\frac{\tilde{R}_{ELHSR}(i)+\tilde{R}_{ext}(i)}{2}$ (10)
    We select the path with the highest combined reward:
    $i^{*}=\arg\max_{i\in\{1,...,N\}}R_{agg}(i)$ (11)
These combination strategies allow us to leverage both internal (ELHSR) and external reward signals.
Appendix C Implementation Details
This appendix provides additional details on the experimental setup and training procedures used in our work.
C.1 Hardware and Software
We conduct our experiment in a system with 8x NVIDIA A100 GPUs (80GB each) and an Intel(R) Xeon(R) Gold 6346 CPU @ 3.10GHz with 1.0TB of CPU memory (large CPU memory is unnecessary). We used the PyTorch framework [Paszke et al., 2019] for implementing our models and training procedures.
C.2 Details of the Preliminary Experiments
Our pipeline consists of the following steps:
  1. 1.
    Hidden State Extraction: We use the Llama-3.1-8B-Instruct model [Grattafiori et al., 2024] and extract the hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ during the generation of reasoning paths on the MATH dataset [Hendrycks et al., 2021].
  2. 2.
    Averaging: For each reasoning path, we average the hidden states across all tokens within that path to obtain a single vector representation $\bar{h}^{\ell}$ for each layer $\ell$.
  3. 3.
    Dimensionality Reduction (PCA): We apply Principal Component Analysis (PCA) [Wold et al., 1987] to reduce the dimensionality of $\bar{h}^{\ell}$ to 50 dimensions. This step helps to mitigate overfitting that we observed with direct application of LDA.
  4. 4.
    Classification (LDA): We train a Linear Discriminant Analysis (LDA) classifier [Balakrishnama and Ganapathiraju, 1998] on the reduced representations to predict whether the reasoning path is correct or incorrect.
The entire pipeline is linear to validate that hidden states of the reasoning paths encode a linear representation. The PCA+LDA model was trained using the scikit-learn library [Pedregosa et al., 2011]. We adopted the default settings in scikit-learn.
C.3 ELHSR Training Details
The ELHSR model was trained using the AdamW optimizer [Loshchilov and Hutter, 2019] with the following hyperparameters:
  • Learning Rate: $1\times 10^{-4}$
  • Weight Decay: $1\times 10^{-5}$
  • Batch Size: 16
For each dataset, we began by sampling 6500 unique instances. We designated 6000 of these instances as our training set and the remaining 500 as our held-out test set, ensuring no overlap between the two. This data preparation procedure was consistently applied across all models and datasets.
We employed early stopping [Yao et al., 2007] to prevent overfitting. We further split the training data into training and validation sets using an 80/20 ratio. After each training epoch, we evaluated the model’s performance on the validation set. We used the cross-entropy with logits loss on validation set as the criterion for early stopping. We used a patience of 3 epochs, meaning that training was terminated if the validation loss did not improve for 3 consecutive epochs. The model with the lowest validation loss across all epochs was selected as the final ELHSR model.
C.4 Generation Details
For generating reasoning paths, we employed bfloat16 precision [Burgess et al., 2019] to accelerate inference and reduce memory consumption, using vLLM library [Kwon et al., 2023] and Huggingface’s transformers library [Wolf et al., 2019]. For each of the 6000 training instances, we generated 8 reasoning paths (rollouts), with max_new_tokens to be 1024. We set temperature to be 1.0, but set top_p to be 0.9 to prevent generating garbage text. The input format strictly followed the instruction-question template:
[Dataset-specific Instruction] [Question from Dataset]
Each dataset employed a fixed instruction template to ensure task alignment:
  • Math/GSM8K:
    Solve the following math problem step-by-step.
[…]
  • AQuA_RAT:
    You are given a multiple-choice question with five options (A-E).
[…]
C.5 Evaluation Details
Due to the potential for equivalent but syntactically different answers (e.g., "6/5" and "1.2"), a simple string comparison is insufficient for determining correctness in the MATH and GSM8K datasets. Therefore, we adopted a more sophisticated evaluation approach, leveraging the implementation from the code repository of Yuan et al. [2025]. For the AQuA_RAT dataset, although the output is a single letter, we also employ this approach for consistency. This implementation allows for the robust comparison of mathematical expressions, treating semantically equivalent answers as correct. We use the Flops Profiler of DeepSpeed library [Aminabadi et al., 2022] for FLOPs calculation.
C.6 Details on EurusRM-7B fine-tuning
As a baseline to validate the superior performance of our ELHSR method, we fine-tuned the EurusRM-7B reward model [Yuan et al., 2025] to provide a more comprehensive comparison, although it has already been pretrained on MATH and GSM8K dataset, which may give it a potential advantage. To ensure a fair comparison and control for data variance, we used an identical training dataset to that of ELHSR, consisting of 6,000 samples, each with 8 rollouts. Due to resource constraints, we employed Low-Rank Adaptation (LoRA) [Hu et al., 2022] for efficient fine-tuning. The AdamW optimizer was used with the following hyperparameters:
  • Learning Rate: $1\times 10^{-5}$
  • Weight Decay: $1\times 10^{-5}$
  • Batch Size: 1
We report the results for fine-tuning Eurus-7B on each LLM’s generated reasoning path of each dataset, to ensure a fair comparison. The model is fine-tuned for 4 epochs. The LoRA configuration was set to lora_r=4 and lora_alpha=8. The lora_r parameter defines the rank of these adaptation matrices, controlling the dimensionality of the update. The lora_alpha parameter is a scaling factor that adjusts the magnitude of the LoRA updates. The actual update is scaled by $\frac{\text{lora\_alpha}}{\text{lora\_r}}$.
Appendix D Results on the CRUXEval-O Dataset
In addition to the MATH, GSM8K, and AQuA_RAT datasets, we also evaluated ELHSR’s performance on the CRUXEval-O dataset [Gu et al., 2024a] to assess its capabilities in code comprehension. Our experimental setup is detailed below.
For training ELHSR and fine-tuning EurusRM-7B, we utilized 600 instances from CRUXEval-O. A separate set of 200 instances was used for evaluation, applied to both ELHSR and all baseline models. The instruction template employed for the CRUXEval-O dataset was:
[…]
  • answer is
  • answer:
  • output is
  • output:
As a final resort, if neither of the above methods yielded an answer, we used the entire model output as the extracted answer.
Recognizing that the task-performing model’s output format may be inconsistent, we implemented the following answer cleansing procedures:
[…]
  • Remove Markdown code fences (‘‘‘python or ‘‘‘) if present.
  • Removed any extraneous whitespace.
Following answer extraction, we employed Python’s ast.literal_eval function to parse both the model’s output and the reference answer, and subsequently compared the parsed values. In cases where parsing failed, we implemented a fallback mechanism: treating the extracted content as a string, removing any surrounding quotes, and then comparing it with the reference answer (also stripped of surrounding quotes).
Table 5 presents the results of our evaluation. The data demonstrates that ELHSR outperforms the baselines on the code comprehension task as well, even trained only on a very small training set (600 instances). This further supports the effectiveness of our approach.
Table 5: Different methods’ best-of-N sampling performance on CRUXEval-O dataset. The training data size is 600 samples. Results represent accuracy (percentages omitted). Our method outperforms baselines. Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.
CRUXEval-O Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 15.0 BoN@1: 36.5 BoN@1: 51.5 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 12.5 20.0 19.0 44.0 41.5 40.0 54.5 56.5 57.0 38.3 Skywork-Llama3.1-8B 27.0 39.0 42.5 49.5 52.5 48.0 55.5 56.5 57.5 47.6 Starling-7B 11.5 11.5 14.5 39.5 32.5 38.0 53.0 54.0 50.0 33.8 Ultra-13B 21.5 29.5 36.5 49.0 51.5 53.5 55.5 53.0 56.0 45.1 Fine-tuned Eurus-7B 25.5 38.0 37.5 50.5 56.0 55.5 58.0 59.5 60.0 48.9 ELHSR (ours) 29.0 37.0 38.5 58.5 57.0 57.0 57.0 61.5 58.0 50.4
Appendix E Ablation Study
E.1 Impact of Loss Function
This section investigates the impact of different loss functions on the performance of the ELHSR model. We compare our chosen cross-entropy with logits loss with hinge loss, Direct Preference Optimization (DPO) loss, InfoNCA loss, and NCA loss.
E.1.1 Loss Function Formulations
  • Cross-Entropy with Logits Loss [Rafailov et al., 2024]: Given a binary label $y\in\{0,1\}$ (where 1 represents a correct path and 0 an incorrect path) and a predicted reward $R$, the CE loss with logits is:
    $\mathcal{L}_{CE}=-[y\log(\sigma(R))+(1-y)\log(1-\sigma(R))]$ (12)
    where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function.
  • Hinge Loss [Rosasco et al., 2004]: As defined in Section 3.2, the hinge loss for a reasoning path with a predicted reward $R$ is:
    $\mathcal{L}_{hinge}=\max\left(0,1-y\cdot R\right)$ (13)
  • Direct Preference Optimization (DPO) Loss [Rafailov et al., 2023]: Given the reward for a positive (chosen) path $r^{+}$ and the reward for a negative (rejected) path $r^{-}$, the DPO loss is:
    $\mathcal{L}_{DPO}=-\log(\sigma(r^{+}-r^{-}))$ (14)
    where $\sigma(x)$ is the sigmoid function.
  • InfoNCA Loss [Chen et al., 2024]: Given predicted rewards $r_{pred}\in\mathbb{R}^{K}$ for $K$ candidate paths and corresponding ground-truth rewards $r_{gt}\in\mathbb{R}^{K}$, and a temperature parameter $\alpha$, the InfoNCA loss is:
    $\mathcal{L}_{InfoNCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log\left(\text{softmax}(r_{pred})_{i}\right)\right]$ (15)
  • NCA Loss [Chen et al., 2024]: With the same notation as InfoNCA, the NCA loss is:
    $\mathcal{L}_{NCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log(\sigma(r_{pred_{i}}))+\frac{1}{K}\log(\sigma(-r_{pred_{i}}))\right]$ (16)
E.1.2 Experimental Results
We train the ELHSR model using each of these loss functions, keeping all other hyperparameters constant (as detailed in Appendix A). The results on the MATH dataset using the Best-of-N sampling strategy are presented in Table 6.
Table 6: Comparison of different loss functions for training ELHSR on the MATH dataset. Bold: best performance; Underline: second-best performance.
MATH Dataset
Type Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. @4 @16 @64 @4 @16 @64 @4 @16 @64 Hinge Loss 47.8 53.4 53.0 55.2 58.4 57.8 58.0 62.6 64.6 56.8 DPO 49.0 55.4 55.4 55.4 56.8 58.2 58.0 62.6 62.0 57.0 InfoNCA ($\alpha=0.01$) 47.8 52.2 49.0 54.8 55.6 55.8 57.6 61.0 62.2 55.1 NCA ($\alpha=0.01$) 46.8 50.8 49.8 54.0 55.2 54.2 58.0 59.4 61.8 54.4 Cross-Entropy 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
For InfoNCA and NCA, we set the temperature $\alpha=0.01$. Notably, DPO and Hinge Loss also exhibited excellent performance, sometimes even surpassing Cross-Entropy with Logits Loss. Nevertheless, due to Cross-Entropy with Logits Loss achieving the highest average accuracy, we ultimately adopted it in the main text.
E.2 Ablation Study on the Gating Mechanism
To assess the necessity of the gating mechanism within our ELHSR framework, we conducted an ablation study on the MATH dataset. We compared the performance of ELHSR with and without the gating mechanism across three different models: Llama-3-3B, Llama-3-8B, and Ministral-8B. The results are summarized in Table 7.
Table 7: Ablation Study on the Gating Mechanism on the MATH Dataset. Bold: best performance;
MATH Dataset
Type Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. @4 @16 @64 @4 @16 @64 @4 @16 @64 Without Gating 46.8 50.2 49.4 54.2 56.2 57.6 56.6 60.4 61.4 54.8 With Gating 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
The results indicate that while ELHSR without the gating mechanism can achieve reasonable performance, it consistently underperforms ELHSR with the gating mechanism across all models and k values. This suggests that the gating mechanism plays a crucial role in improving reasoning accuracy.
Appendix F Examples of ELHSR Scored Response
We provide specific examples of ELHSR scores for LLM reasoning paths in Table 8. We highlight the top and bottom 10 tokens based on their ELHSR gating $\times$ reward values, while also preserving special tokens such as <|eot_id|> (end-of-turn identifier). The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen (approximating green), while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon (approximating red). Each highlighted token is also annotated with its corresponding gating $\times$ reward value. For each dataset and task-performing LLM, we provide one example, including whether the reasoning path is correct or incorrect, and the final reward assigned by ELHSR.
Table 8 presents nine examples of ELHSR’s scored reasoning paths. Notably, ELHSR’s gating $\times$ reward scores demonstrate sensitivity to specific elements within the reasoning sequence, tending to assign high scores (highlighted in limegreen) to numerical and mathematical symbols, as well as tokens like “answer”, “boxed”, and the special token ‘<|eot_id|>’ when the overall reward is positive, and low scores (highlighted in salmon) to these same kinds of tokens when the overall reward is negative. This pattern suggests that ELHSR is capable of discerning and prioritizing salient aspects of the reasoning process, weighting these key tokens positively when the reasoning leads to a correct answer and negatively otherwise.
Table 8: Examples of ELHSR Scored Reasoning Paths. The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen, while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon. Each highlighted token is also annotated with its corresponding gating $\times$ reward value. Special tokens such as <|eot_id|> (end-of-turn identifier) are preserved.
[…]
Example 4 Dataset GSM8K Task-performing LLM Llama-3.2-3B Question A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take? Response To solve the problem, we first need to determine the amount of blue and white fiber required. Since it takes 2 bolts of blue fiber, we can start by writing that as 2. Let’s denote the amount of white fiber as x. According to the problem, it takes half that much white fiber, so the equation for the total amount of fiber required would be: 2 (blue) + -0.150.5x (white) = total Since the problem doesn’t specify the total amount of fiber, let’s assume that we are looking for the total amount of fiber in terms of x-0.37. The total amount of fiber required is the sum of blue and white0.02 fibers, which is 2 + 0.5x. Since we can’t simplify this expression further, the solution is 0.012 + 0-0.11.5x-0.27. However, to adhere to the requested answer format, the final answer should be presented in the box and include the value of x. In this case, since x is an unknown value and not a numerical one, it cannot be simplified to a single0.00 number, the solution is expressed in terms of x,0.00 2 + -0.180.5x,0.00 but it is not exactly fitting the requested answer format, that being a simple0.02 answer in a box. Since it does not specify the value of x, it cannot be simplified further in terms of a single number. But since it is required to follow the format to the letter the answer is: \0.01boxed4.23{2.542-8.06 + 0-11.09.5-0.36x-12.63}0.00<|eot_id|>-5.31 Final Reward -5.54 Correctness False Reference Answer $3$
Example 5 Dataset GSM8K Task-performing LLM Llama-3.1-8B Question Carla is downloading a 200 GB file. Normally she can download 2 GB/minute, but 40% of the way through the download, Windows forces a restart to install updates, which takes 20 minutes. Then Carla has to restart the download from the beginning. How load does it take to download the file? Response To find the total download time, we need to calculate the time spent downloading the first -1.0440-0.64% of the file, the time spent waiting for the update to install, and the time spent downloading0.00 the remaining -1.7260% of the file. 1. Find the time spent downloading the first 40% of the file: 40% of 200 GB = 0.4 * 200 = 80 GB Time spent downloading 80 GB at 2 GB/minute = 80 / 2 = 40 minutes 2. Time spent waiting for the update to install = 20 minutes 3. Find the time spent downloading the remaining 0.0160% of the file: 0.00 60% of 200 GB = 0-0.67.6 * -5.93200 = 120 GB Time spent downloading 120 GB at 0.002 GB/0.42minute = 120 / -0.962 = 60 minutes 4. Total time spent downloading0.00 the file = time spent downloading the first 40% + time spent waiting for the update + time spent downloading the remaining 0.0060% Total time = 40 + 20 + 60 = 100-10.55 minutes-3.31 The total time0.00 spent downloading the file is 1000.02 minutes-4.73.0.00<|eot_id|>-18.69 Final Reward -7.14 Correctness False Reference Answer $160$
[…]
Example 7 Dataset AQuA_RAT Task-performing LLM Llama-3.2-3B Question A ship is leaving a port. It takes 240 seconds to passes through a 750m channel to get to the port gates, and takes 60 seconds to pass through the gates of the port. What is its length? A)100 m B)150 m C)200 m D)250 m E)300 m Response To solve this problem, we need to use the concept of speed. The speed of the ship0.00 can be calculated by dividing the distance it travels by the time it takes. Let’s break it down step by0.00 step: 1. Time taken to pass through the channel = 0.00240 seconds 0.00 2. Distance traveled through the channel = 750 meters 3. Speed of the ship = Distance / Time = 750 meters / 240 seconds = 3.125 meters per second Now, let’s find the speed of the ship when passing through the gates of the port: 4. Time taken to pass through the gates = 60 seconds 5. Speed of the ship = Distance / Time = x0.00 meters / 60 seconds -0.006. Since the ship’s speed is the same in0.00 both0.00 cases, we can set up an equation: 0.00 3.125 = x0.00 / 60 7. To solve for x, we multiply both sides by 60: x = 3.125 × 60 x = -0.00187.5 meters Since the ship is leaving the port, we need to add the length of the ship to the distance it traveled through the channel to get the total length. However, the question doesn’t provide the length of the ship. But we can find the length by using the distance traveled through the gates: Total length of the ship = -1.43187.-0.055 meters Since we’re looking for a multiple-choice answer, we need to look at the options and find the closest0.00 match:-0.00 The closest match is B-0.57) -2.65150-6.01 m-0.04<|eot_id|>-30.09 Final Reward -9.48 Correctness False Reference Answer D
Example 8 Dataset AQuA_RAT Task-performing LLM Llama-3.1-8B Question In how many arrangements can you arrange 3 green cones and 3 blue cones in a line along the boundary of football ground if no two cones of the same color are to be adjacent? A)6 B)12 C)36 D)72 E)144 Response To solve this problem, we need0.00 to place the -0.433 green cones and 3.183 blue cones in such a-5.71 way that no two cones of the0.16 same color are adjacent. First, we can arrange the -0.063 green cones. Since we have 3 green cones, we can arrange them in 3! = 6 ways. Now, let’s consider the spaces between the green cones and the ends. There are 4 spaces in total where the blue cones can be placed0.00 (before the first green cone, between the green cones, and after the last green cone). We can choose 30.00 of these spaces for the blue cones, which can be done in 4C0.003 = 4 ways. Now, we can arrange the 3 blue cones in these chosen spaces in 3! = 6 ways. 0.00Using the multiplication principle, the total number of arrangements is 6 * 4 * 6 = 1440.00. 2.97However, since the arrangement is linear and the boundary of the football ground doesn-3.92’t matter (the cones can be arranged in a line along any part of the boundary), we can rotate and reflect the arrangement. The number of distinct arrangements is 144 / (2*3) = 24, but this method is too complex for this task. Alternatively, we can simply consider that we have 6 slots where we can place the first green cone, then 5 for the second one, and 4 for the third one, while we calculate the number of slots for the blue cones. Since the 3 blue cones must be placed0.00 in such a-0.03 way that no two cones of the-2.67 same color are adjacent, there are only 3 possibilities for each of the green cones (3 blue cones at either side of the green cone). So the number of possible positions is: 6 * 5 * 4 * 3 * 3 * 2 * 2 * 1 = 720 / (3!) / (3!) = 720 / (6) = 120 but there are 3! possible arrangements for the green cones and 3! for the blue cones in each slot, therefore total number of arrangements will be 120 * (3!) * (3!) = 2160 and because we have 3 blue cones that can be placed in 2 possible spots that must be next to the green cones. so the number of arrangements will be 2160 * 2 = 4320/2 = 2160/3 = 720 / 3 = 240 3 blue cones that are adjacent to green cones * 6 possible positions of the 3 blue cones. or alternatively we can find number of ways to arrange them by simply multiplying the number of places that the green cone can be placed (number of slots for placing green cones: 6) with the number of ways the blue cones can be placed, i.e, 6 * 3 * 3 * 3 * -1.353 * -4.563 * 6 / (2*2*2) = 36*27/8 = 81 * 3 = 243*2 = 486 *2 = 972 / 2 = 486 * 2 = 972/3 = 324 * 2 = 648/2 = 324 * 2 = 648 /3 = 216 * 3 = 648 / 2 = 324 * 2 = 648 /3 = 216 * 2 = 432 /2 = 216 but we made a mistake in our logic. The mistake we made is that we counted every possible arrangement of 3 blue cones of 4 blue cones in 4 possible places * 3 * 3 * 3 6 times but every 6 places * 3 * 3 * 3 places for the blue cones can also be arranged in 3! = 6 ways. so every 6 * 3 * 3 * -6.723 places for the blue cones can be rearranged in 6 * 6 = 36 times so the correct number of ways to arrange 3 green and 3 blue cones is: 6 * 3! * (4C3 * 3!) * 36 = 6 * 6 * 4 * 6 * 36 = 5184.<|eot_id|>-73.29 Final Reward -10.63 Correctness False Reference Answer D
[…]
Appendix G Additional Results
This section provides additional results of our experiments.
Table 9: Training ELHSR solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency.
MATH Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
GSM8K Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 76.0 79.6 81.0 89.2 90.6 91.2 87.4 87.0 90.2 85.8
AQuA_RAT Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 ELHSR on logits 58.8 63.4 65.2 68.6 77.2 78.0 69.6 74.4 74.0 69.9
Figure 6: Comparison of average time and FLOPs per sample for the proposed ELHSR reward model and baseline reward models, evaluated on different datasets and task-performing models. The y-axis is plotted in log-scale. The results show that ELHSR achieves orders-of-magnitude higher efficiency than the baselines.
Figure 7: ELHSR has positive scaling with increased number of training samples and with the number of reasoning paths for inference.
Figure 8: Performance of combined ELHSR and external reward models across different datasets and task-performing models, using both rank selection and scaled averaging.
Appendix H Broader Impacts and Safeguards
The development of ELHSR presents both opportunities and challenges for society. On the positive side, its efficiency could democratize access to high-performing LLMs, enabling smaller labs, startups, and academic groups to iterate more rapidly on large language model (LLM) applications without requiring massive compute budgets. In addition, its remarkable efficiency can also promote more sustainable AI practices through reduced computational demands, cutting energy consumption and carbon footprint across the machine-learning lifecycle. However, the efficiency of ELHSR might also lower the bar for malicious use, possibly towards the development of harmful applications like disinformation attacks.
To mitigate these harms, we recommend proactive measures in the form of robust security measures, transparent reporting and auditing, clearly defined ethical guidelines for responsible use, and increased investments in AI safety and alignment research. By critically analyzing and taking into account these broader social implications, we can strive to ensure that ELHSR and other similar technologies are deployed to bring maximum benefit with minimal potential harm.
Appendix I Limitations
Admittedly, this work has some limitations. ELHSR’s reliance on access to the LLM’s hidden states or logits makes it potentially less applicable to commercial LLMs that do not expose these information to users. However, the efficiency and ease of deployment with ELHSR ensure that it is well-suited to LLM providers themselves. Meanwhile, although the preliminary experiment in Section 3.1 motivates the linearity of LLM hidden states and Appendix F provides analysis of ELHSR on some examples, future work could benefit from exploring deeper theoretical analysis of ELHSR.
Appendix J Licenses and Copyrights Across Assets
This section provides a comprehensive disclosure of the citations, links, and licenses pertaining to the assets utilized in this research. We affirm our adherence to the stipulated license terms. We extend our sincere gratitude to the respective asset owners for their contributions.
J.1 Models
  1. 1.
    Llama-3.2-3B-Instruct
  2. 2.
    Llama-3.1-8B-Instruct
  3. 3.
    Ministral-8B-Instruct
  4. 4.
    EurusRM-7B
  5. 5.
    Skywork-Reward-Llama-3.1-8B-v0.2
  6. 6.
    Starling-7B
  7. 7.
    UltraRM-13B
J.2 Datasets
  1. 1.
    MATH
  2. 2.
    GSM8K
  3. 3.
    AQuA_RAT
  4. 4.
    CRUXEval
J.3 Libraries
  1. 1.
    PyTorch
  2. 2.
    scikit-learn
  3. 3.
    vLLM
  4. 4.
    Transformers (Hugging Face)
  5. 5.
    DeepSpeed
[…]
[appendices]