### v3

# Mining Intrinsic Rewards from LLM Hidden States for Efficient Best-of-N Sampling

Abstract.
Best-of-N sampling is a powerful method for improving Large Language Model (LLM) performance, but it is often limited by its dependence on massive, text-based reward models. These models are not only computationally expensive but also data-hungry, requiring extensive labeled datasets for training. This creates a significant data challenge, as they overlook a rich, readily available data source: the LLM’s own internal hidden states. To address this data and efficiency gap, we introduce SWIFT (Simple Weighted Intrinsic Feedback Technique), a novel and lightweight method that learns a reward function directly from the rich information embedded in LLM hidden states. Operating at the token embedding level, SWIFT employs simple linear layers to effectively distinguish between preferred and dispreferred generations, eliminating the need for computationally intensive text-based modeling. Extensive experiments on standard benchmarks show that SWIFT outperforms existing baselines (12.7% higher accuracy than EurusRM-7B on MATH dataset) while using less than 0.005% of their parameters. Its robust scalability, compatibility with certain closed-source models via logit access, and ability to combine with traditional reward models for additional performance highlight SWIFT’s practical value and contribution to more efficient data-driven LLM post-training. Our code is available at https://github.com/aster2024/SWIFT.
Large Language Models, Knowledge Discovery, Reward Modeling
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780302isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processingccs: Computing methodologies Learning from implicit feedbackccs: Computing methodologies Learning latent representations
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18140886.
1. Introduction
Figure 1. An illustration comparing traditional reward model and SWIFT.
An illustration comparing traditional reward model and SWIFT.
The proliferation of Large Language Models (LLMs) has unlocked unprecedented capabilities in various domains, from complex reasoning to user assistance (Liu et al., 2023; Zhang et al., 2023; Samel et al., 2023). To further refine and enhance these capabilities after initial pretraining, various post-training techniques are employed, including methods that scale computation at test-time to improve performance on difficult tasks (Liu et al., 2025). A key one of these test-time scaling strategies is Best-of-N sampling, in which a reward model selects the best response from a set of candidates (Snell et al., 2025). However, the effectiveness of this approach is significantly constrained by the reward models themselves. Current methods rely on large, independently trained neural networks that predict rewards based on generated text. These models are not only computationally expensive but also data-hungry, requiring massive datasets and incurring substantial overhead during both training and inference (Guo et al., 2025; Namgoong et al., 2024; Zhang et al., 2024). This exclusive reliance on textual data necessitates complex architectures to capture subtle errors, posing a major barrier to efficient LLM enhancement.
We argue that this paradigm overlooks a rich, readily available data source: the task-performing LLM’s own internal hidden states. During generation, an LLM naturally emits internal signals that reflect its confidence of its answer and the quality of its reasoning. Seminal work has shown that this “internal knowledge” is not only information-rich but is often captured in a linearly separable representations within the model’s hidden states (Zhang et al., 2025; Xie et al., 2024). Our own preliminary experiments support this finding (Section 4.1), revealing a clear linear relationship between hidden states and reasoning correctness. This property is significant, as it points to a path for highly efficient reward modeling: rather than collecting ever-larger text datasets to train ever-larger external models, we can mine the intrinsic signals already present within the LLM itself.
However, leveraging this insight introduces a fundamental architectural challenge. Traditional reward models are designed to process discrete text tokens, creating a mismatch with the continuous, high-dimensional vectors of hidden states. Addressing this mismatch requires a new architecture tailored specifically to extract meaningful signals from these hidden representations. While prior work has explored hidden states to improve reasoning, such methods typically focus on single, final-answer tokens or rely on subjective definitions of reasoning steps (Burns et al., 2023; Azaria and Mitchell, 2023). As a result, they offer limited performance gains. They also do not produce a direct reward signal suitable for Best-of-N sampling.
To address this challenge, we propose the Simple Weighted Intrinsic Feedback Technique (SWIFT), a novel architecture specifically designed to learn from hidden states at the token level (Figure 1). For each token, SWIFT learns both a linear gating value and a linear reward projection. The final reward is computed as a weighted average of these token-level rewards, offering a holistic, fine-grained assessment of the entire generation. Our experiments show that SWIFT outperforms baseline reward models (demonstrated 12.7% higher accuracy in average on MATH dataset than EurusRM-7B, a state-of-the-art open-source reward model on math and code domain (Yuan et al., 2025)) while using fewer than 0.005% of their parameters. It is exceptionally data-efficient, can be trained with only a handful of examples, and achieves orders-of-magnitude improvements in computational efficiency (FLOPs) compared to existing baselines.
Our code is available at https://github.com/aster2024/SWIFT, and the project website is on https://aster2024.github.io/swift-website/. In summary, our main contributions are:
  • We introduce SWIFT, a token-level reward model that directly mines LLM hidden states. We demonstrate SWIFT’s orthogonal design can augment traditional reward models to achieve further performance gains.
  • We extend SWIFT to operate with logit-only access, demonstrating its practical applicability to a broader range of models, including certain closed-source LLMs.
  • Our experiments demonstrate that SWIFT is highly data-efficient, achieving strong performance with minimal training data. Furthermore, it exhibits robust scalability with performance improvement from both training-time and test-time scaling.
2. Related Work
Reward Modeling for Language Models’ Reasoning
The rapid advancement of Large Language Models (LLMs) in complex tasks, such as answering user’s questions (Shen et al., 2023; Mo et al., 2023) and reasoning (Zhou et al., 2024; Xu et al., 2024), has highlighted the pivotal role of reward modeling in both training and post-training inference (Cui et al., 2024; Lambert et al., 2025). Current reward models typically build upon an LLM backbone, with substantial parameter count and require significant amount of data for training, which introduces high computational costs (Guo et al., 2025; Namgoong et al., 2024). Although prior work has proposed various approaches to mitigate such issue (Chan et al., 2024; Fisch et al., 2024), achieving a satisfactory balance between efficiency and performance remains a challenge. This work introduces a novel perspective by leveraging internal representations to achieve a significant improvement in efficiency with satisfactory performance.
Understanding and Leveraging the Internal Representations of Language Models
The rapid development of LLMs has spurred considerable research into their internal representations (Ferrando et al., 2024; Li et al., 2023). Hidden states of LLMs have been effectively utilized in diverse fields, including safety alignment (Zou et al., 2023; Barez et al., 2025), faithfulness and hallucination detection (Burns et al., 2023), knowledge editing (Meng et al., 2022; Yu et al., 2024), and knowledge distillation or transfer (Dasgupta and Cohn, 2025; Wu et al., 2024). By contrast, we effectively leverage the hidden states in intermediate layers of LLMs as a direct signal for reward modeling in reasoning tasks.
3. Problem Statement
A common approach to enhance LLM reasoning is best-of-N sampling (Snell et al., 2025). Let $\mathcal{V}$ be the token vocabulary and let $D$ be a distribution over problem-answer pairs $(x,y)\in\mathcal{V}^{n}\times\mathcal{Y}$, where $y$ is the hidden reference solution for question $x$. Conditioned on $x$, the base LLM $\pi$ draws $N$ reasoning paths independently (i.i.d. draws from $\pi(\cdot\mid x)$), producing $r_{i}=(r_{i,1},\dots,r_{i,m_{i}})\in\mathcal{V}^{m_{i}},\quad i=1,\dots,N,$ where $m_{i}$ is the random length of the $i$-th path. We write $\mathcal{V}^{*}=\bigcup_{\ell=0}^{\infty}\mathcal{V}^{\ell}$. Simultaneously, $\pi$ emits hidden states $h_{i}=(h_{i,1},\dots,h_{i,m_{i}})\;\in\;(\mathbb{R}^{L\times d})^{m_{i}},$ where $L$ is the number of transformer layers, $d$ is the per-layer hidden-state dimension, and each $h_{i,j}\in\mathbb{R}^{L\times d}$ is the stacked hidden state after emitting $r_{i,j}$. Define the correctness indicator $F:\mathcal{V}^{*}\times\mathcal{Y}\to\{0,1\},\text{s.t.}\,F(r,y)=1\iff r\text{ yields reference answer }y.$ We write $(\mathbb{R}^{L\times d})^{*}=\bigcup_{k=0}^{\infty}(\mathbb{R}^{L\times d})^{k}$. A parametric reward model $R_{\theta}:(\mathbb{R}^{L\times d})^{*}\to\mathbb{R}$ assigns a scalar score to each $h_{i}$ (hence for each $r_{i}$). Note that neither $\pi$ nor $R_{\theta}$ sees $y$. At test time we select $i^{*}=\arg\max_{1\leq i\leq N}R_{\theta}(h_{i})$ and return $r_{i^{*}}$. Our goal is to choose $\theta$ to maximize the probability that $r_{i^{*}}$ is correct (expectation of $F(r_{i^{*}},y)$) under $D$:
(1) $\max_{\theta}\;\mathbb{E}_{(x,y)\sim D}\;\mathbb{E}_{(r_{1},\dots,r_{N})\sim\pi(\cdot\mid x)}\bigl[F\bigl(r_{\arg\max_{i}R_{\theta}(h_{i})},y\bigr)\bigr].$
4. Methods
4.1. Motivation
Figure 2. Cross-validation accuracy of the PCA+LDA pipeline for predicting correctness at each layer of Llama-3.1-8B. The results demonstrate that hidden states contain information about reasoning correctness with linear representation.
Cross-validation accuracy of the PCA+LDA pipeline for predicting correctness at each layer of Llama-3.1-8B.
We first present a preliminary experiment demonstrating the potential for enhancing reasoning using hidden states. Recent work has shown that LLM hidden states contain information about the correctness probability of a given reasoning path, often exhibiting linear representations (Zhang et al., 2025; Xie et al., 2024). To verify this, we designed a simple pipeline to predict the correctness of reasoning steps based on the LLM’s hidden states.
Our pipeline extracts hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ from the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) during reasoning on the MATH dataset (Hendrycks et al., 2021), averages the hidden states across all tokens within each reasoning path to obtain a single vector representation $\bar{h}^{\ell}$ for each layer $\ell$, then reduces $\bar{h}^{\ell}$ for each $\ell$ to 50 dimensions using Principal Component Analysis (PCA) (Wold et al., 1987), and classifies the reasoning as correct or incorrect using Linear Discriminant Analysis (LDA) (Balakrishnama and Ganapathiraju, 1998). The pipeline is linear to verify that the hidden states can encode a linear representation for confidence of the answer. For more details, please refer to Appendix D.2.
We performed 5-fold cross-validation on 3000 instances of the MATH dataset. For each layer $\ell$, we trained the pipeline on the training folds and evaluated the classification accuracy on the held-out test fold. Figure 2 presents the average cross-validation accuracy for each layer.
As shown in Figure 2, the accuracy is approximately 80% on each layer, indicating a strong signal for reasoning correctness within the hidden states. This result verifies the aforementioned theory and motivates the development of a novel, lightweight, linear reward model that directly leverages this information, as described in the following sections.
4.2. Our Approach - SWIFT
Therefore, building upon the observation that LLM internal representations exhibit linear properties, we designed SWIFT (Simple Weighted Intrinsic Feedback Technique) model to be a parameter-efficient, linear model, which learns a token-level reward function directly from the concatenated hidden states.
For each token $t$ in a reasoning path, we concatenate and flatten the hidden states (after residual connection) from all $L$ layers: $h_{t}=[h^{1}_{t};h^{2}_{t};...;h^{L}_{t}]\in\mathbb{R}^{Ld}$, where $d$ is the dimension of the hidden state. We then apply a linear transformation to obtain a gating value $g_{t}$ and a token-level reward $r_{t}$:
(2) $\begin{pmatrix}\tilde{g}_{t}\\ r_{t}\end{pmatrix}=W_{SWIFT}h_{t}+b_{SWIFT}$
where $W_{SWIFT}\in\mathbb{R}^{2\times Ld}$ is the weight matrix, $b_{SWIFT}\in\mathbb{R}^{2}$ is the bias vector, $\tilde{g}_{t}$ is the pre-activation gating value, and $r_{t}$ is the token-level reward. The gating value is then passed through a sigmoid function (Han and Moraga, 1995):
(3) $g_{t}=\sigma(\tilde{g}_{t})=\frac{1}{1+e^{-\tilde{g}_{t}}}$
The final reward $R$ for the entire reasoning path is the weighted average of the token-level rewards, using the gating values as weights:
(4) $R=\frac{\sum_{t=1}^{T}(g_{t}\cdot r_{t})}{\max\left(\sum_{t=1}^{T}g_{t},\epsilon\right)}$
where $T$ is the number of tokens, and $\epsilon$ is a small constant (e.g., $10^{-8}$) for numerical stability. This model only contains $O(L\times d)$ parameters, which is significantly fewer than those in traditional reward models.
Since different tokens contribute differently to the overall correctness, the gating values allow the model to assign higher weights to more important tokens. In addition, we conduct ablation experiment to show that the gating mechanism can improve performance in Appendix E.2.
We use the binary cross-entropy with logits loss (Rafailov et al., 2024) for training. Let $y\in\{0,1\}$ denote the correctness label of a reasoning path, where $y=1$ indicates a correct path and $y=0$ indicates an incorrect path. For a path with hidden state representation $h$, the model predicts a scalar logit score $R(h)\in\mathbb{R}$. The binary cross-entropy with logits loss is:
(5) $\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\cdot\log(\sigma(R(h_{i})))+(1-y_{i})\cdot\log(1-\sigma(R(h_{i})))\right]$
where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function.
Although binary cross-entropy with logits loss generally yields good results, our experiments detailed in Appendix E.1 demonstrate that DPO (Rafailov et al., 2023) and hinge loss (Rosasco et al., 2004) can also perform well, sometimes even surpassing cross-entropy. However, given that cross-entropy performs best on average, we adopt it as our primary loss function in the main text. A comparison study among several loss functions is provided in Appendix E.1.
5. Experiments
5.1. Experimental Setup
Our experiments are conducted across mathematical reasoning, code comprehension, commonsense abilities and symbolic reasoning tasks. We evaluate our method on the MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), AQuA_RAT (Ling et al., 2017), Imbue Code Comprehension (Imbue Team, 2024), HellaSwag (Zellers et al., 2019) and CoinFlip (Wei et al., 2022) datasets. The models used for generation are Llama-3.2-3B-Instruct (Meta, 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Ministral-8B-Instruct (Mistral AI, 2024). The generation temperature is 1.0 with top_p to be 0.9. We use 6000 instances (each with 8 reasoning paths) for training and 500 instances for evaluation. When training SWIFT, we use early stopping (Yao et al., 2007) based on validation set performance to mitigate overfitting. For more details on the generation process, training and evaluation, please refer to Appendix D.3, D.4 and D.5.
We compare our method against six recent open-source reward models: EurusRM-7B (Yuan et al., 2025), Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024), Starling-7B (Zhu et al., 2024), UltraRM-13B (Cui et al., 2024), RLHFlow-8B-Deepseek (Xiong et al., 2024), Math-Shepherd-7B (Wang et al., 2024). These baselines are very parameter heavy and trained on extensive datasets, while achieving leading performance among open-source reward models. These six baseline reward models have all been explicitly trained on data related to reasoning and code domains. Their enhanced performance in these specific areas is well-documented and validated in their respective original publications. It is particularly noteworthy that the development of Eurus-7B, Ultra-13B, RLHFlow-8B-Deepseek, and Math-Shepherd-7B involved training and evaluation on the MATH and GSM8K datasets. This corresponds directly with the datasets used in our work, confirming the relevance of these models as strong baselines.
For links and brief descriptions of baselines, please refer to Appendix B. For parameter count and training data size of baselines compared against SWIFT, see Table 3 for details.
To provide a more comprehensive comparison, we also evaluated a fine-tuned version of EurusRM-7B in Table 12, with details illustrated in Appendix D.6, specifically fine-tuned on the relevant datasets and model-generated solutions used in our experiments, while using the same loss function as in the original publication to ensure consistency. This is despite EurusRM-7B already being pre-trained on MATH and GSM8K, which gives it a potential advantage.
5.2. Accuracy
Table 1. Different methods’ best-of-N sampling performance on mathematics domain, including MATH, GSM8K, and AQuA_RAT datasets. Results represent accuracy (percentages omitted). Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.
Different methods’ best-of-N sampling performance on mathematics domain, including MATH, GSM8K, and AQuA_RAT datasets.
MATH Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 42.6 48.2 46.8 50.8 52.0 52.2 54.6 56.8 55.0 51.0 Skywork-Llama3.1-8B (Liu et al., 2024) 43.8 48.4 48.8 52.2 52.4 53.4 56.8 59.0 61.6 52.9 Starling-7B (Zhu et al., 2024) 39.6 41.2 39.8 50.4 49.0 49.0 53.8 50.2 47.0 46.7 Ultra-13B (Cui et al., 2024) 44.6 47.4 44.4 53.0 50.6 50.4 53.6 53.0 54.0 50.1 RLHFlow-8B-Deepseek (Xiong et al., 2024) 43.2 46.4 47.6 51.2 49.6 49.8 56.2 58.4 57.8 51.1 Math-Shepherd-7B (Wang et al., 2024) 43.8 42.6 43.6 48.8 50.8 49.0 55.8 58.8 54.8 49.8 Fine-tuned Eurus-7B 42.2 47.0 46.6 51.8 53.0 50.2 54.8 57.2 57.2 51.1 SWIFT (ours) 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
GSM8K Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 53.0 BoN@1: 80.4 BoN@1: 79.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 69.8 76.2 82.4 87.0 89.0 90.4 87.0 90.8 90.4 84.8 Skywork-Llama3.1-8B (Liu et al., 2024) 77.2 81.8 85.2 89.6 89.4 89.2 83.0 85.8 87.4 85.4 Starling-7B (Zhu et al., 2024) 62.4 65.6 71.4 83.0 87.2 85.4 78.4 81.2 80.0 77.2 Ultra-13B (Cui et al., 2024) 72.6 78.6 82.4 88.0 88.6 86.8 82.6 84.6 82.0 82.9 RLHFlow-8B-Deepseek (Xiong et al., 2024) 52.4 53.8 60.0 80.4 77.0 80.0 88.4 91.2 93.4 75.2 Math-Shepherd-7B (Wang et al., 2024) 54.8 49.8 43.4 80.6 75.4 73.0 84.8 86.8 86.8 70.6 Fine-tuned Eurus-7B 79.2 84.4 86.8 90.0 91.2 91.0 88.2 91.8 92.8 88.4 SWIFT (ours) 80.0 86.4 87.6 90.6 89.6 90.6 89.8 93.2 93.4 89.0
AQuA_RAT Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 42.0 BoN@1: 53.6 BoN@1: 56.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 44.6 42.8 44.6 58.6 60.6 56.8 48.4 34.6 27.2 46.5 Skywork-Llama3.1-8B (Liu et al., 2024) 60.4 68.0 69.4 69.0 77.0 79.2 65.6 71.2 72.0 70.2 Starling-7B (Zhu et al., 2024) 51.4 55.0 56.6 64.2 70.8 69.8 56.4 48.6 42.8 57.3 Ultra-13B (Cui et al., 2024) 58.4 61.6 64.8 67.4 72.8 71.8 58.4 51.4 48.0 61.6 RLHFlow-8B-Deepseek (Xiong et al., 2024) 53.8 54.0 48.8 63.6 65.0 61.4 64.6 63.8 62.6 59.7 Math-Shepherd-7B (Wang et al., 2024) 52.2 53.4 53.0 64.0 67.8 66.2 63.6 60.4 60.2 60.1 Fine-tuned Eurus-7B 57.4 66.6 62.0 68.8 72.8 69.8 71.2 71.8 71.0 67.9 SWIFT (ours) 61.0 70.4 70.8 69.6 77.0 77.0 75.2 75.8 78.0 72.8
Table 2. Different methods’ best-of-N sampling performance on general domain, including Imbue Code Comprehension, HellaSwag and CoinFlip datasets. Results represent accuracy (percentages omitted). Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.
Different methods’ best-of-N sampling performance on general domain, including Imbue Code Comprehension, HellaSwag and CoinFlip datasets.
Imbue Code Comprehension Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 17.2 BoN@1: 47.6 BoN@1: 54.4 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 26.0 28.0 30.2 60.0 62.2 66.2 59.0 61.6 64.0 50.8 Skywork-Llama3.1-8B (Liu et al., 2024) 32.6 47.4 56.0 67.0 70.8 75.2 66.6 71.4 74.8 62.4 Starling-7B (Zhu et al., 2024) 18.8 21.0 21.0 57.6 62.6 64.2 58.6 60.4 62.6 47.4 Ultra-13B (Cui et al., 2024) 28.0 34.8 38.8 61.8 66.6 68.6 60.6 64.4 65.2 54.3 RLHFlow-8B-Deepseek (Xiong et al., 2024) 29.2 37.0 44.4 66.4 71.4 72.0 68.4 74.0 75.2 59.8 Math-Shepherd-7B (Wang et al., 2024) 22.6 25.4 25.4 60.6 60.0 55.8 63.0 65.4 61.4 48.8 Fine-tuned Eurus-7B 35.6 54.4 60.4 67.0 69.4 76.0 67.8 71.0 71.8 63.7 SWIFT (ours) 36.8 57.6 65.6 74.0 84.0 85.0 74.6 82.4 85.4 71.7
HellaSwag Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 55.2 BoN@1: 61.2 BoN@1: 68.2 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 67.4 71.6 72.0 72.6 73.6 74.0 77.8 77.4 78.0 73.8 Skywork-Llama3.1-8B (Liu et al., 2024) 63.0 64.0 67.0 70.6 71.6 71.8 73.0 74.0 76.4 70.2 Starling-7B (Zhu et al., 2024) 53.6 56.2 59.0 64.8 65.8 63.6 66.0 65.0 63.6 62.0 Ultra-13B (Cui et al., 2024) 60.8 64.0 64.2 69.6 69.4 68.2 67.8 67.8 70.6 66.9 RLHFlow-8B-Deepseek (Xiong et al., 2024) 65.2 68.4 69.2 68.2 70.8 71.0 75.2 78.2 79.8 71.8 Math-Shepherd-7B (Wang et al., 2024) 59.8 61.0 62.2 66.0 65.2 63.8 70.2 69.8 73.0 65.7 Fine-tuned Eurus-7B 70.4 76.2 77.4 75.2 79.4 78.6 80.4 81.4 81.6 77.8 SWIFT (ours) 71.0 75.8 77.4 78.0 82.4 83.2 83.4 88.0 88.2 80.8
CoinFlip Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 56.2 BoN@1: 62.6 BoN@1: 20.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 65.6 64.2 64.6 83.6 85.6 91.8 45.8 69.8 77.4 72.0 Skywork-Llama3.1-8B (Liu et al., 2024) 72.6 81.0 88.0 88.0 94.4 97.4 25.2 28.2 26.0 66.8 Starling-7B (Zhu et al., 2024) 66.6 73.0 76.2 79.0 84.2 87.2 22.0 19.6 21.0 58.8 Ultra-13B (Cui et al., 2024) 61.6 62.0 59.4 82.0 85.6 84.0 14.0 11.4 11.4 52.4 RLHFlow-8B-Deepseek (Xiong et al., 2024) 64.0 65.6 75.6 83.4 96.4 99.4 34.6 49.4 72.8 71.2 Math-Shepherd-7B (Wang et al., 2024) 55.8 46.4 35.4 62.4 40.4 15.2 20.0 19.0 15.4 34.4 Fine-tuned Eurus-7B 91.2 99.4 99.8 95.2 99.8 100.0 57.8 95.8 99.8 93.2 SWIFT (ours) 91.6 99.8 100.0 95.4 99.6 100.0 57.8 96.2 99.6 93.3
Table 1 reports the Best-of-N performance on mathematics domain of each reward model on the MATH, GSM8K, and AQuA_RAT datasets, and Table 2 reports their performance on other general domains, including code comprehension (Imbue Code Comprehension dataset), commonsense (HellaSwag dataset) and symbolic reasoning (CoinFlip dataset). Our SWIFT model outperforms all baseline reward models across most BoN settings. For example, SWIFT is 12.7% more accurate on average than EurusRM-7B on the MATH dataset, with EurusRM-7B being a state-of-the-art open-source reward model on the math and code domains (Yuan et al., 2025). A notable observation is that while many baseline reward models perform well on specific subsets of the data, SWIFT demonstrates consistent and robust performance across the entire dataset. Although the fine-tuned EurusRM-7B achieves respectable results, it still falls short of SWIFT in accuracy and incurs substantially higher computational costs during both training and inference. In contrast, SWIFT not only achieves the best overall performance but does so with a parameter footprint of less than 0.005% of the baselines and is trained on only 6,000 samples per dataset (see Section 5.3 for detailed comparison). This highlights the exceptional efficiency of our approach.
It is noteworthy that the results presented here were achieved without extensive hyperparameter tuning. Further performance gains could likely be realized through more refined optimization strategies. For example, exploring alternative loss functions, as demonstrated in Appendix Table 7, increasing the training dataset size based on the scalability analysis in Section 5.6, or training SWIFT on a subset of layers as explored in Section 5.7 (Table 5), could yield further improvements. This suggests that SWIFT has significant potential for further optimization and improvement.
5.3. Efficiency Analysis
Figure 3. Comparison of average time and FLOPs per sample for SWIFT and baselines, averaged across different datasets and task-performing models with 1-sigma error bars using standard error of the mean. The y-axis is plotted in log-scale. It shows that SWIFT achieves orders-of-magnitude higher efficiency than baselines. See Appendix Figure 6 and Figure 7 for full details.
Comparison of average time and FLOPs per sample for SWIFT and baselines, averaged across different datasets and task-performing models with 1-sigma error bars using standard error of the mean. The y-axis is plotted in log-scale.
We demonstrate SWIFT’s significant parameter efficiency and data efficiency in Table 3. Moreover, we evaluate the computational efficiency of SWIFT by comparing the average time and FLOPs required to assign a reward to each sample. Figure 3 presents a comparative analysis of SWIFT and baseline reward models across various datasets and task-performing models. The results show that SWIFT achieves orders-of-magnitude improvements in both time and FLOPs compared to the baselines. Notably, as shown in Section 5.6, even a significantly smaller training dataset leads to satisfactory performance. This substantial reduction in computational cost, combined with its parameter efficiency and data efficiency, makes SWIFT a highly practical solution for reward modeling in both resource-constrained environments and large-scale applications.
[…]
Table 3. Comparison of Reward Model Parameters and Training Data Size (number of questions $\times$ number of answers). Our method is remarkably efficient in terms of parameters and data.
Different methods’ best-of-N sampling performance on MATH, GSM8K, and AQuA_RAT datasets.
Reward Model Parameters Training data size
Eurus-7B $7.1\times 10^{9}$ $5.8\times 10^{5}\times 2$
Skywork-Llama3.1-8B $7.5\times 10^{9}$ $8.0\times 10^{4}\times 2$
Starling-7B $6.7\times 10^{9}$ $1.8\times 10^{5}\times 7$
Ultra-13B $1.3\times 10^{10}$ $7.5\times 10^{5}\times 2$
RLHFlow-8B-Deepseek $8.0\times 10^{9}$ $2.5\times 10^{5}\times 1$
Math-Shepherd-7B $7.2\times 10^{9}$ $4.5\times 10^{5}\times 1$
SWIFT (on Llama-3.2-3B) $1.8\times 10^{5}$ $6.0\times 10^{3}\times 8$
SWIFT (on Llama-3.1-8B) $2.7\times 10^{5}$ $6.0\times 10^{3}\times 8$
SWIFT (on Ministral-8B) $3.0\times 10^{5}$ $6.0\times 10^{3}\times 8$
Figure 4. SWIFT has positive scaling with increased number of training samples and with the number of reasoning paths for inference. The result is averaged across datasets and task-performing models.
SWIFT has positive scaling with increased number of training samples and with the number of reasoning paths for inference.
5.4. Generalization Test
To assess transferability, we pretrain SWIFT on 10k samples of DeepScaleR dataset (Luo et al., 2025) and evaluate it directly on downstream benchmarks (no dataset-specific fine-tuning). With Ministral-8B and BoN@64, SWIFT achieves 62.8/93.6/75.8 on MATH/GSM8K/AQuA_RAT, consistently exceeding all baselines (e.g., fine-tuned Eurus: 57.2/92.8/71.0; Skywork: 61.6/87.4/72.0) and achieving performance comparable to fine-tuned SWIFT. In addition, to further validate that the strong performance of SWIFT does not benefit from fine-tuning solely, we fine-tune Skywork-Llama3.1-8B (the strongest baseline RM on MATH) and evaluate BoN@64 on MATH. Across task LLMs Llama3-3B/Llama3-8B/Ministral-8B, fine-tuned Skywork yields 52.2/54.0/62.6, while SWIFT achieves 53.6/62.6/62.8, consistently outperforming the fine-tuned baseline. This demonstrates that SWIFT’s improvements stem from its design rather than fine-tuning alone, and confirms robust cross-dataset generalization.
5.5. Domains beyond reasoning accuracy (helpfulness/safety)
Table 4. Alignment performance on the PKU-SafeRLHF dataset evaluated by GPT-4o. We report BoN@4 accuracy using Ministral-8B as the policy. SWIFT consistently outperforms strong baselines on both helpfulness and safety dimensions.
Dimension SWIFT Eurus Skywork Ultra RLHFlow Starling Shepherd
Helpfulness 64.2 59.8 63.2 54.6 61.0 60.4 53.0
Safety 92.2 89.8 91.6 79.0 90.2 91.6 84.2
Finally, to test whether SWIFT extends to alignment-oriented evaluation, we run Best-of-$N$ on PKU-SafeRLHF dataset (Ji et al., 2025) using GPT-4o as judge (Ministral-8B; BoN@4). The result is characterized in Table 4. SWIFT significantly outperforms all large text-based reward models. The consistent helpfulness and safety improvement over these baselines suggest that SWIFT’s hidden-state representations capture nuanced response quality beyond reasoning correctness, making it suitable for general alignment tasks.
5.6. Scalability with Training Set Size and Reasoning Paths
Scaling large language models (LLMs) can be achieved through various strategies, broadly categorized as training-time scaling and test-time scaling (Snell et al., 2025; Kaplan et al., 2020). In this section, we investigate the scalability of SWIFT with respect to both training set size and the number of reasoning paths used during inference.
From Figure 4, our results demonstrate that SWIFT’s accuracy consistently improves with both the number of training samples and the size of the reasoning paths set. Specifically, under different BoN@k settings, we find that SWIFT’s accuracy increases monotonically with training set size. Likewise, for any given training set size, SWIFT’s accuracy also improves as the number of reasoning paths used in the BoN sampling process increases.
It is interesting to note that, as shown in Figure 4, the performance of SWIFT with training-time scaling does not appear to plateau even with 6000 training samples, which is the one used in the experiments of Section 5.2. This suggests that further increasing the training set size may yield additional performance improvements, underscoring the scalability of SWIFT. These results highlight SWIFT as a compelling and practical approach to scaling reward modeling for real-world applications.
5.7. Maintaining Performance with Selective Layer Utilization for Higher Efficiency
Table 5. SWIFT’s high Best-of-N sampling performance is maintained even when trained only on layers 16, 24, 28, and 32 for Llama-3.1-8B, leading to a substantial reduction in parameters and improved efficiency.
Method MATH GSM8K AQuA_RAT Avg.
BoN@1: 47.2 BoN@1: 80.4 BoN@1: 53.6
@4 @16 @64 @4 @16 @64 @4 @16 @64
SWIFT on partial layers 55.8 59.0 60.0 89.4 90.8 92.0 69.6 76.8 77.4 74.5
SWIFT on all layers 55.4 59.4 62.6 90.6 89.6 90.6 69.6 77.0 77.0 74.6
While SWIFT leverages hidden states from all layers of the LLM, we investigated the potential for further parameter reduction by selectively using hidden states from only a subset of layers. Specifically, for the Llama-3.1-8B model, which comprises 32 layers (plus an embedding layer), we explored training SWIFT using only the hidden states from layers 16, 24, 28, and 32, following the setting in Azaria and Mitchell (2023). This configuration reduces the parameter count to approximately 4/33 of the full-layer SWIFT model. The results, presented in Table 5, demonstrate that performance remains comparable to the full-layer SWIFT, with certain settings even showing improved accuracy. Notably, on the GSM8K dataset, the BoN@16 and BoN@64 accuracies surpass those of the full-layer SWIFT. As shown in Table 1, the performance of the layer-selected SWIFT also consistently exceeds that of baselines with significantly larger parameter counts, further underscoring the efficiency and effectiveness of our approach.
We further find that later layers encode stronger correctness signals: using only the first 4 layers gives BoN@64 accuracy 38.2/44.6/45.4 (Llama3-3B/Llama3-8B/Ministral-8B on MATH), while using only the last 4 layers improves it to 55.4/60.4/61.8. Please refer to Appendix E.4 for details.
5.8. Logit-Based Training for Closed-Source LLMs
Table 6. Training SWIFT solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency. For results on GSM8K and AQuA_RAT dataset, please refer to Appendix Table 12.
MATH Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
While SWIFT offers significant advantages in terms of efficiency and performance, making it well-suited for deployment by large model providers and end-users, a limitation arises when working with closed-source LLMs where hidden states are inaccessible. To overcome this constraint, we leverage the fact that some closed-source models still provide access to their logits, such as GPT-3.5-turbo (OpenAI, 2023a) and GPT-4 (Achiam et al., 2023), as demonstrated by Finlayson et al. (2024) and OpenAI (2023b). We train SWIFT directly on these logits, resulting in a model with a parameter count that is orders of magnitude smaller than traditional reward models - and even smaller than SWIFT trained on all layers. This enables seamless deployment on personal devices such as smartphones and personal computers. As demonstrated in Table 6 and Appendix Table 12, the logit-based SWIFT maintains strong performance. Furthermore, as shown in Table 1, its performance even surpasses that of many baselines. These results highlight the versatility and adaptability of SWIFT, making it a viable solution even when access to hidden states is restricted.
Concretely, with BoN@64, logit-only SWIFT reaches 50.6/55.8/57.4 on MATH, 81.0/91.2/90.2 on GSM8K, and 65.2/78.0/74.0 on AQuA_RAT for Llama-3.2-3B/Llama-3.1-8B/Ministral-8B, respectively, demonstrating that informative correctness signals can be extracted even from logit access alone.
5.9. Combining SWIFT with External Reward Models
Figure 5. Average performance of combined SWIFT and external reward models across different datasets and task-performing models, for both rank selection and scaled averaging. See Appendix Figure 8 and Figure 9 for full details.
Average performance of combined SWIFT and external reward models across different datasets and task-performing models.
SWIFT offers a highly parameter-efficient approach to reward modeling, enabling its seamless integration with existing external reward models. Given SWIFT’s minimal parameter footprint (less than 0.005% of baseline models), the additional computational overhead introduced by incorporating SWIFT is negligible compared to relying solely on the external reward model. We explore two straightforward methods for combining SWIFT with external rewards: rank selection and scaled averaging. In rank selection, we choose the sample with the best rank as determined by either SWIFT or the external reward model. In scaled averaging, we compute the average of the rewards from both models after normalizing them to the [0, 1] range. Detailed descriptions of these methods are provided in Appendix C.
We evaluated the effectiveness of these combination strategies by integrating SWIFT with various traditional reward models across different datasets and task-performing models, assessing performance using best-of-N sampling. Figure 5 presents the average results across different datasets for both ranking and scaling methods, applied to various task-performing models. Detailed results for each dataset and task-performing model can be found in Appendix Figure 8 and Figure 9. Our findings indicate that combining SWIFT with external reward models can lead to further improvements in accuracy, particularly on the GSM8K and AQuA-RAT datasets. The MATH dataset did not exhibit significant performance gains, potentially because SWIFT already significantly outperforms the external reward models on this benchmark. This performance boost is likely due to SWIFT and external reward models capturing different types of errors, resulting in improved overall accuracy-particularly with a higher number of best-of-N reasoning paths. This combination strategy offers a pathway to further performance gains without incurring significant computational overhead.
5.10. Analysis on SWIFT
To investigate the mechanism of SWIFT, we present illustrative examples of its scoring of reasoning paths in Appendix F. The examples with the highest and lowest gating $\times$ reward values are highlighted, along with their corresponding numerical scores. Notably, SWIFT assigns relatively high or low gating $\times$ reward values to specific numerical and mathematical symbols, as well as particular tokens such as “answer”, “boxed” and the special token ‘<|eot_id|>’. This behavior suggests that SWIFT possesses the ability to identify and weigh key components within the reasoning process.
In addition, we also conduct ablation study on the gating mechanism of SWIFT. SWIFT includes a lightweight token-level gating module that re-weights token representations before aggregating them into a path-level score. We ablate this component by removing gating while keeping the same training budget and evaluation protocol. On MATH, gating yields consistent gains across all BoN settings. For BoN@64, accuracies improve from 49.4 to 53.6 (Llama-3.2-3B), 57.6 to 62.6 (Llama-3.1-8B), and 61.4 to 62.8 (Ministral-8B). Similar trends hold at BoN@4 and BoN@16, confirming that gating is beneficial across varying numbers of candidate paths. This token-level selection mechanism allows SWIFT to emphasize semantically decisive portions of reasoning traces, such as intermediate steps, numerical computations, and conclusion markers, while downweighting boilerplate text. The details can be found in Table 8.
To better understand where SWIFT’s gains come from, we further break down improvements by MATH sub-domain (Ministral-8B as the task LLM; BoN@64). SWIFT improves over Math-Shepherd-7B across all sub-domains, with gains ranging from +2.44 points (Geometry) to +10.52 points (Counting & Probability), and consistently strong gains on Algebra (+8.87), Intermediate Algebra (+10.31), and Number Theory (+9.68). See Table 9 for the full breakdown.
We also compare alternative objectives for training SWIFT on MATH, including hinge-style and preference-style losses. Across three task LLMs (Llama3-3B/Llama3-8B/Ministral-8B) and three BoN settings (@4, @16, @64), cross-entropy achieves the best average accuracy (57.5), followed by DPO (57.0), hinge loss (56.8), InfoNCA (55.1) and NCA (54.4). The relatively small gap between the top three suggests that SWIFT is robust to the choice of training signal. Cross-entropy remains our default due to its consistent edge and simpler implementation. The detailed experiment is illustrated in Appendix E.1.
6. Conclusion
In this paper, we presented SWIFT, a novel reward modeling approach for LLMs that leverages the rich information contained in their internal hidden states. SWIFT offers both high efficiency and effectiveness, and it scales well during both training and inference. We further showed that comparable performance can be achieved by training SWIFT on a subset of layers or even on logits alone, significantly reducing computational costs and making it more versatile. The successful integration of SWIFT with standard reward models underscores its flexibility and potential.
Admittedly, this work has some limitations. SWIFT’s reliance on access to the LLM’s hidden states or logits may reduce its applicability to commercial LLMs that do not expose this information to users. However, its efficiency and ease of deployment make SWIFT especially well-suited to LLM providers themselves. In addition, SWIFT must be trained separately for each LLM as hidden-state dimensionality and layer count vary across models. Nonetheless, given its small model size and minimal data requirements, the associated computational overhead is negligible, especially when compared to the cost of training the task-performing LLM itself.
Acknowledgements.
This work was supported by the National Science Foundation (NSF) under Grant No. III-2106758. We gratefully acknowledge Mrs. Jiahe Jin from Zhiyuan College, Shanghai Jiao Tong University, for her invaluable assistance in creating the initial version of our illustrations. We also thank Mr. Zhihui Xie from The University of Hong Kong for offering precious suggestions.
Appendix Contents
Appendix A Relevance to KDD Community
Our work is highly relevant to the KDD community and aligns directly with the topic of Modern AI and Big Data of the KDD 2026 Research Track, since according to the description of the call for paper page in the official website, deep representation learning, generation and reasoning are all contained in this topic. We address a critical challenge in the large-scale deployment of Large Language Models (LLMs) by shifting the focus of the data source for reward modeling. Instead of relying on computationally prohibitive, text-based reward models, our work introduces SWIFT, a technique that mines an LLM’s internal hidden states to extract an intrinsic signal of reasoning correctness. This approach embodies the principles of deep representation learning by leveraging the rich information encoded within a model’s internal layers. Furthermore, it directly contributes to advancements in generation and reasoning by offering a lightweight and highly efficient method to enhance output quality. By demonstrating that valuable, actionable knowledge can be discovered from a model’s internal states, which is a novel data source, our paper presents a significant contribution to the efficient and scalable application of large language models (LLMs), a core topic of interest to the KDD community.
Appendix B Baseline Reward Models Description
To provide further context for the baseline reward models used in our experiments, we present a brief overview and links to their respective Hugging Face Model Hub pages:
  • EurusRM-7B: EurusRM-7B is a 7B parameter reward model trained on a mixture of UltraInteract (Yuan et al., 2025), UltraFeedback (Cui et al., 2024), and UltraSafety datasets (Guo et al., 2024), with a focus on improving reasoning performance. EurusRM-7B stands out as the best 7B RM overall and achieves similar or better performance than much larger baselines. Particularly, it outperforms GPT-4 in certain tasks.
  • Skywork-Reward-Llama-3.1-8B-v0.2: This 8B parameter reward model is built on the Llama-3 architecture (Grattafiori et al., 2024) and trained on a curated dataset of 80K high-quality preference pairs. It excels at handling preferences in complex scenarios, including challenging preference pairs, and span various domains such as mathematics, coding, and safety. As of October 2024, Skywork-Reward-Llama-3.1-8B-v0.2 ranks first among 8B models.
  • Starling-RM-7B: Starling-RM-7B is a reward model trained from Llama2-7B-Chat (Touvron et al., 2023), following the method of training reward models in the instructGPT paper (Ouyang et al., 2022). It is trained with the Nectar preference dataset (Zhu et al., 2024), which is based on GPT-4’s preferences.
  • UltraRM-13B: UltraRM is a reward model initialized by Llama2-13B (Touvron et al., 2023) and fine-tuned on the UltraFeedback dataset (Cui et al., 2024). It achieves state-of-the-art performance among open-source reward models on public preference test sets.
  • RLHFlowRM-8B-Deepseek: RLHFlowRM-8B-Deepseek is trained from Llama-3.1-8B-Instruct on Deepseek generated data. It is evaluated on GSM8K and MATH dataset and demonstrated significant performance gains.
  • Math-Shepherd-7B: Math-Shepherd is trained with automatic process annotation. It is benchmarked on GSM8K and MATH dataset and substantially boosted the task-performing LLM’s performance.
Appendix C Detailed Approach for Combining SWIFT with External Reward Models
In the Best-of-N (BoN) setting, we explore two strategies for combining the SWIFT reward ($R_{SWIFT}$) with an external reward signal ($R_{ext}$), typically from a traditional reward model. We have $N$ reasoning paths.
  1. (1)
    Ranking-based Combination: Let $rank_{SWIFT}(i)$ and
    $rank_{ext}(i)$ be the ranks of the $i$-th reasoning path according to SWIFT and the external reward model, respectively (lower rank is better). The relative rank is:
    (6) $rank_{rel}(i)=\max(rank_{SWIFT}(i),rank_{ext}(i))$
    We select the path with the lowest relative rank:
    (7) $i^{*}=\arg\min_{i\in\{1,...,N\}}rank_{rel}(i)$
  2. (2)
    Scaling-based Combination: We linearly scale both rewards to the range [0, 1] for the $N$ paths [^1] If all the rewards are equal, we set them all to $1.0$, but this is almost impossible in practice.:
    (8) $\tilde{R}_{SWIFT}(i)=\frac{R_{SWIFT}(i)-\min_{j}R_{SWIFT}(j)}{\max_{j}R_{SWIFT}(j)-\min_{j}R_{SWIFT}(j)}$
    (9) $\tilde{R}_{ext}(i)=\frac{R_{ext}(i)-\min_{j}R_{ext}(j)}{\max_{j}R_{ext}(j)-\min_{j}R_{ext}(j)}$
    The aggregated reward is:
    (10) $R_{agg}(i)=\frac{\tilde{R}_{SWIFT}(i)+\tilde{R}_{ext}(i)}{2}$
    We select the path with the highest combined reward:
    (11) $i^{*}=\arg\max_{i\in\{1,...,N\}}R_{agg}(i)$
These combination strategies allow us to leverage both internal (SWIFT) and external reward signals.
Appendix D Implementation Details
This appendix provides additional details on the experimental setup and training procedures used in our work.
D.1. Hardware and Software
We conduct our experiment in a system with 8x NVIDIA A100 GPUs (80GB each) and an Intel(R) Xeon(R) Gold 6346 CPU @ 3.10GHz with 1.0TB of CPU memory (large CPU memory is unnecessary). We used the PyTorch framework (Paszke et al., 2019) for implementing our models and training procedures.
D.2. Details of the Preliminary Experiments
Our pipeline consists of the following steps:
  1. (1)
    Hidden State Extraction: We use the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) and extract the hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ during the generation of reasoning paths on the MATH dataset (Hendrycks et al., 2021).
  2. (2)
    Averaging: For each reasoning path, we average the hidden states across all tokens within that path to obtain a single vector representation $\bar{h}^{\ell}$ for each layer $\ell$.
  3. (3)
    Dimensionality Reduction (PCA): We apply Principal Component Analysis (PCA) (Wold et al., 1987) to reduce the dimensionality of $\bar{h}^{\ell}$ to 50 dimensions. This step helps to mitigate overfitting that we observed with direct application of LDA.
  4. (4)
    Classification (LDA): We train a Linear Discriminant Analysis (LDA) classifier (Balakrishnama and Ganapathiraju, 1998) on the reduced representations to predict whether the reasoning path is correct or incorrect.
The entire pipeline is linear to validate that hidden states of the reasoning paths encode a linear representation. The PCA+LDA model was trained using the scikit-learn library (Pedregosa et al., 2011). We adopted the default settings in scikit-learn.
D.3. SWIFT Training Details
The SWIFT model was trained using the AdamW optimizer (Loshchilov and Hutter, 2019) with the following hyperparameters:
  • Learning Rate: $1\times 10^{-4}$
  • Weight Decay: $1\times 10^{-5}$
  • Batch Size: 16
For each dataset, we created a training set of 6000 instances, drawn from the original training set, and a held-out test set of 500 instances, drawn from the original validation set (due to the absence of labels in some datasets’ test sets). This consistent data preparation approach was applied across all datasets, with the exceptions of GSM8K (where the original test set was used due to the lack of a validation set) and Imbue Code Comprehension (where we performed a manual train/test split as no split was provided).
In addition, we need to generate reasoning path and determine whether each reasoning path is correct. The generation details are provided in Appendix D.4 and the method of determining each reasoning path’s correctness is detailed in Appendix D.5.
We employed early stopping (Yao et al., 2007) to prevent overfitting. We further split the training data into training and validation sets using an 80/20 ratio. After each training epoch, we evaluated the model’s performance on the validation set. We used the cross-entropy with logits loss on validation set as the criterion for early stopping. We used a patience of 3 epochs, meaning that training was terminated if the validation loss did not improve for 3 consecutive epochs. The model with the lowest validation loss across all epochs was selected as the final SWIFT model.
D.4. Generation Details
For generating reasoning paths, we employed bfloat16 precision (Burgess et al., 2019) to accelerate inference and reduce memory consumption, using vLLM library (Kwon et al., 2023) and Huggingface’s transformers library (Wolf et al., 2019). For each of the 6000 training instances, we generated 8 reasoning paths (rollouts), with max_new_tokens to be 1024. We set temperature to be 1.0, but set top_p to be 0.9 to prevent generating garbage text. The input format strictly followed the instruction-question template:
[Dataset-specific Instruction] [Question from Dataset]
Each dataset employed a fixed instruction template to ensure task alignment:
  • Math/GSM8K:
    Solve the following math problem step-by-step.
[…]
  • AQuA_RAT:
    You are given a multiple-choice question with five options (A-E).
[…]
  • Imbue Code Comprehension:
    Analyze the following problem step-by-step. The question includes a list of choices.
[…]
  • HellaSwag:
    You are given a context and four possible continuations (A--D). Decide which option best continues the context.
[…]
  • CoinFlip:
    Answer the following Boolean question with detailed reasoning.
[…]
D.5. Evaluation Details / Method to identify correctness
This section details how we determine each reasoning path’s correctness for both training and evaluation, and the method for FLOPs calculation.
Due to the potential for equivalent but syntactically different answers (e.g., "6/5" and "1.2"), a simple string comparison is insufficient for determining correctness in the MATH and GSM8K datasets. Therefore, we adopted a more sophisticated evaluation approach, leveraging the implementation from the code repository of Yuan et al. (2025). For the AQuA_RAT, HellaSwag and CoinFlip dataset, although the output is a single letter, we also employ this approach for consistency. This implementation allows for the robust comparison of mathematical expressions, treating semantically equivalent answers as correct.
[…]
  • answer is
  • answer:
  • output is
  • output:
As a final resort, if neither of the above methods yielded an answer, we used the entire model output as the extracted answer.
Recognizing that the task-performing model’s output format may be inconsistent, we implemented the following answer cleansing procedures:
  • Removed any occurrences of $\text{…}$‘ by replacing them with the inner text.
  • Replaced escaped braces $\{$‘ and ’\}‘ with regular braces ’{‘ and ’}‘.
  • Remove Markdown code fences (‘‘‘python or ‘‘‘) if present.
  • Removed any extraneous whitespace.
Following answer extraction, we employed Python’s
ast.literal_eval function to parse both the model’s output and the reference answer, and subsequently compared the parsed values. In cases where parsing failed, we implemented a fallback mechanism: treating the extracted content as a string, removing any surrounding quotes, and then comparing it with the reference answer (also stripped of surrounding quotes).
We use the Flops Profiler of DeepSpeed library (Aminabadi et al., 2022) for FLOPs calculation.
D.6. Details on EurusRM-7B fine-tuning
As a baseline to validate the superior performance of our SWIFT method, we fine-tuned the EurusRM-7B reward model (Yuan et al., 2025) to provide a more comprehensive comparison, although it has already been pretrained on MATH and GSM8K dataset, which may give it a potential advantage. To ensure a fair comparison and control for data variance, we used an identical training dataset to that of SWIFT, consisting of 6,000 samples, each with 8 rollouts. Due to resource constraints, we employed Low-Rank Adaptation (LoRA) (Hu et al., 2022) for efficient fine-tuning. The AdamW optimizer was used with the following hyperparameters:
  • Learning Rate: $1\times 10^{-5}$
  • Weight Decay: $1\times 10^{-5}$
  • Batch Size: 1
Our fine-tuning process employs a loss function identical to that of the original paper Yuan et al. (2025). The specific loss function, $\mathcal{L}_{\text{ULTRAINTERACT}}$, is defined as follows:
[…]
In this formulation, $x$ denotes the problem prompt, $y_{c}$ represents the chosen (correct) answer, and $y_{r}$ is the rejected (incorrect) answer. The reward model, parameterized by $\theta$, is denoted by $r_{\theta}(\cdot)$, and $\sigma$ is the sigmoid function. The $\mathcal{L}_{\text{BT}}$ term is a Bradley-Terry-style loss that optimizes the relative rewards, encouraging the reward of the correct answer to be higher than that of the incorrect one. The $\mathcal{L}_{\text{DR}}$ term directly regularizes the rewards by increasing the reward for the chosen answer and decreasing it for the rejected answer.
Furthermore, our data sampling methodology is analogous to the approach described in the original paper. For each problem instance, a training pair is constructed by randomly matching one correct solution with one incorrect solution from the same problem.
We report the results for fine-tuning Eurus-7B on each LLM’s generated reasoning path of each dataset, to ensure a fair comparison. The model is fine-tuned for 4 epochs. The LoRA configuration was set to lora_r=4 and lora_alpha=8. The lora_r parameter defines the rank of these adaptation matrices, controlling the dimensionality of the update. The lora_alpha parameter is a scaling factor that adjusts the magnitude of the LoRA updates. The actual update is scaled by $\frac{\text{lora\_alpha}}{\text{lora\_r}}$.
Appendix E Ablation Study
E.1. Impact of Loss Function
This section investigates the impact of different loss functions on the performance of the SWIFT model. We compare our chosen cross-entropy with logits loss with hinge loss, Direct Preference Optimization (DPO) loss, InfoNCA loss, and NCA loss.
E.1.1. Loss Function Formulations
  • Cross-Entropy with Logits Loss (Rafailov et al., 2024): Given a binary label $y\in\{0,1\}$ (where 1 represents a correct path and 0 an incorrect path) and a predicted reward $R$, the CE loss with logits is:
    (13) $\mathcal{L}_{CE}=-[y\log(\sigma(R))+(1-y)\log(1-\sigma(R))]$
    where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function.
  • Hinge Loss (Rosasco et al., 2004): As defined in Section 4.2, the hinge loss for a reasoning path with a predicted reward $R$ is:
    (14) $\mathcal{L}_{hinge}=\max\left(0,1-y\cdot R\right)$
  • Direct Preference Optimization (DPO) Loss (Rafailov et al., 2023): Given the reward for a positive (chosen) path $r^{+}$ and the reward for a negative (rejected) path $r^{-}$, the DPO loss is:
    (15) $\mathcal{L}_{DPO}=-\log(\sigma(r^{+}-r^{-}))$
    where $\sigma(x)$ is the sigmoid function.
  • InfoNCA Loss (Chen et al., 2024): Given predicted rewards $r_{pred}\in\mathbb{R}^{K}$ for $K$ candidate paths and corresponding ground-truth rewards $r_{gt}\in\mathbb{R}^{K}$, and a temperature parameter $\alpha$, the InfoNCA loss is:
    (16) $\mathcal{L}_{InfoNCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log\left(\text{softmax}(r_{pred})_{i}\right)\right]$
  • NCA Loss (Chen et al., 2024): With the same notation as InfoNCA, the NCA loss is:
    (17) $\mathcal{L}_{NCA}=-\sum_{i=1}^{K}\left[\text{softmax}\left(\frac{r_{gt}}{\alpha}\right)_{i}\cdot\log(\sigma(r_{pred_{i}}))+\frac{1}{K}\log(\sigma(-r_{pred_{i}}))\right]$
E.1.2. Experimental Results
We train the SWIFT model using each of these loss functions, keeping all other hyperparameters constant (as detailed in Appendix A). The results on the MATH dataset using the Best-of-N sampling strategy are presented in Table 7.
Table 7. Comparison of different loss functions for training SWIFT on the MATH dataset. Bold: best performance; Underline: second-best performance.
MATH Dataset
Type Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. @4 @16 @64 @4 @16 @64 @4 @16 @64 Hinge Loss 47.8 53.4 53.0 55.2 58.4 57.8 58.0 62.6 64.6 56.8 DPO 49.0 55.4 55.4 55.4 56.8 58.2 58.0 62.6 62.0 57.0 InfoNCA ($\alpha=0.01$) 47.8 52.2 49.0 54.8 55.6 55.8 57.6 61.0 62.2 55.1 NCA ($\alpha=0.01$) 46.8 50.8 49.8 54.0 55.2 54.2 58.0 59.4 61.8 54.4 Cross-Entropy 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
For InfoNCA and NCA, we set the temperature $\alpha=0.01$. Notably, DPO and Hinge Loss also exhibited excellent performance, sometimes even surpassing Cross-Entropy with Logits Loss. Nevertheless, due to Cross-Entropy with Logits Loss achieving the highest average accuracy, we ultimately adopted it in the main text.
E.2. Ablation Study on the Gating Mechanism
To assess the necessity of the gating mechanism within our SWIFT framework, we conducted an ablation study on the MATH dataset. We compared the performance of SWIFT with and without the gating mechanism across three different models: Llama-3-3B, Llama-3-8B, and Ministral-8B. The results are summarized in Table 8.
Table 8. Ablation Study on the Gating Mechanism on the MATH Dataset. Bold: best performance;
MATH Dataset
Type Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. @4 @16 @64 @4 @16 @64 @4 @16 @64 Without Gating 46.8 50.2 49.4 54.2 56.2 57.6 56.6 60.4 61.4 54.8 With Gating 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
The results indicate that while SWIFT without the gating mechanism can achieve reasonable performance, it consistently underperforms SWIFT with the gating mechanism across all models and k values. This suggests that the gating mechanism plays a crucial role in improving reasoning accuracy.
E.3. MATH Sub-domain Breakdown
To better understand when SWIFT helps, we compare SWIFT against Math-Shepherd-7B on MATH and break down the BoN@64 gains by sub-domain (Ministral-8B as the task LLM). Table 9 shows improvements across all sub-domains, with larger gains on categories that typically require more multi-step reasoning.
Table 9. SWIFT gains over Math-Shepherd-7B by MATH sub-domain (Ministral-8B, BoN@64; accuracy difference in percentage points).
Category Gain
Precalculus 7.14
Algebra 8.87
Intermediate Algebra 10.31
Number Theory 9.68
Prealgebra 4.88
Geometry 2.44
Counting & Probability 10.52
Per-category accuracy gains of SWIFT over Math-Shepherd-7B on MATH.
E.4. Layer-wise Ablation Study
Table 10 presents a detailed breakdown of SWIFT’s performance when trained using different subsets of hidden layers from the task-performing LLM. We compare three configurations: using only the first 4 layers, using only the last 4 layers, and using all layers (default). Results are reported for three task LLMs (Llama3-3B, Llama3-8B, Ministral-8B) under BoN@64 on MATH. The first 4 layers yield substantially lower accuracy (38.2/44.6/45.4), confirming that early-layer representations lack sufficient semantic abstraction for reasoning evaluation. In contrast, the last 4 layers achieve 55.4/60.4/61.8, similar to the full model’s performance (53.6/62.6/62.8). This demonstrates that discriminative information for reasoning quality is concentrated in the final layers, allowing SWIFT to remain highly effective even with a reduced layer subset, a useful insight for further efficiency optimization.
Table 10. Layer-wise ablation on MATH (BoN@64). Accuracy (%) across three task LLMs.
Layer Subset Llama3-3B Llama3-8B Ministral-8B
First 4 layers 38.2 44.6 45.4
Last 4 layers 55.4 60.4 61.8
All layers (default) 53.6 62.6 62.8
Appendix F Examples of SWIFT Scored Response
We provide specific examples of SWIFT scores for LLM reasoning paths in Table 11. We highlight the top and bottom 10 tokens based on their SWIFT gating $\times$ reward values, while also preserving special tokens such as <|eot_id|> (end-of-turn identifier). The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen (approximating green), while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon (approximating red). Each highlighted token is also annotated with its corresponding gating $\times$ reward value. For each dataset and task-performing LLM, we provide one example, including whether the reasoning path is correct or incorrect, and the final reward assigned by SWIFT.
Table 11 presents nine examples of SWIFT’s scored reasoning paths. Notably, SWIFT’s gating $\times$ reward scores demonstrate sensitivity to specific elements within the reasoning sequence, tending to assign high scores (highlighted in limegreen) to numerical and mathematical symbols, as well as tokens like “answer”, “boxed”, and the special token ‘<|eot_id|>’ when the overall reward is positive, and low scores (highlighted in salmon) to these same kinds of tokens when the overall reward is negative. This pattern suggests that SWIFT is capable of discerning and prioritizing salient aspects of the reasoning process, weighting these key tokens positively when the reasoning leads to a correct answer and negatively otherwise.
Table 11. Examples of SWIFT Scored Reasoning Paths. The 10 tokens with the highest gating $\times$ reward are highlighted in limegreen, while the 10 tokens with the lowest gating $\times$ reward are highlighted in salmon. Each highlighted token is also annotated with its corresponding gating $\times$ reward value. Special tokens such as <|eot_id|> (end-of-turn identifier) are preserved.
[…]
Example 4 Dataset GSM8K Task-performing LLM Llama-3.2-3B Question A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take? Response To solve the problem, we first need to determine the amount of blue and white fiber required. Since it takes 2 bolts of blue fiber, we can start by writing that as 2. Let’s denote the amount of white fiber as x. According to the problem, it takes half that much white fiber, so the equation for the total amount of fiber required would be: 2 (blue) + -0.150.5x (white) = total Since the problem doesn’t specify the total amount of fiber, let’s assume that we are looking for the total amount of fiber in terms of x-0.37. The total amount of fiber required is the sum of blue and white0.02 fibers, which is 2 + 0.5x. Since we can’t simplify this expression further, the solution is 0.012 + 0-0.11.5x-0.27. However, to adhere to the requested answer format, the final answer should be presented in the box and include the value of x. In this case, since x is an unknown value and not a numerical one, it cannot be simplified to a single0.00 number, the solution is expressed in terms of x,0.00 2 + -0.180.5x,0.00 but it is not exactly fitting the requested answer format, that being a simple0.02 answer in a box. Since it does not specify the value of x, it cannot be simplified further in terms of a single number. But since it is required to follow the format to the letter the answer is: \0.01boxed4.23{2.542-8.06 + 0-11.09.5-0.36x-12.63}0.00<|eot_id|>-5.31 Final Reward -5.54 Correctness False Reference Answer 33
Example 5 Dataset GSM8K Task-performing LLM Llama-3.1-8B Question Carla is downloading a 200 GB file. Normally she can download 2 GB/minute, but 40% of the way through the download, Windows forces a restart to install updates, which takes 20 minutes. Then Carla has to restart the download from the beginning. How load does it take to download the file? Response To find the total download time, we need to calculate the time spent downloading the first -1.0440-0.64% of the file, the time spent waiting for the update to install, and the time spent downloading0.00 the remaining -1.7260% of the file. 1. Find the time spent downloading the first 40% of the file: 40% of 200 GB = 0.4 * 200 = 80 GB Time spent downloading 80 GB at 2 GB/minute = 80 / 2 = 40 minutes 2. Time spent waiting for the update to install = 20 minutes 3. Find the time spent downloading the remaining 0.0160% of the file: 0.00 60% of 200 GB = 0-0.67.6 * -5.93200 = 120 GB Time spent downloading 120 GB at 0.002 GB/0.42minute = 120 / -0.962 = 60 minutes 4. Total time spent downloading0.00 the file = time spent downloading the first 40% + time spent waiting for the update + time spent downloading the remaining 0.0060% Total time = 40 + 20 + 60 = 100-10.55 minutes-3.31 The total time0.00 spent downloading the file is 1000.02 minutes-4.73.0.00<|eot_id|>-18.69 Final Reward -7.14 Correctness False Reference Answer $160$
[…]
Example 7 Dataset AQuA_RAT Task-performing LLM Llama-3.2-3B Question A ship is leaving a port. It takes 240 seconds to passes through a 750m channel to get to the port gates, and takes 60 seconds to pass through the gates of the port. What is its length? A)100 m B)150 m C)200 m D)250 m E)300 m Response To solve this problem, we need to use the concept of speed. The speed of the ship0.00 can be calculated by dividing the distance it travels by the time it takes. Let’s break it down step by0.00 step: 1. Time taken to pass through the channel = 0.00240 seconds 0.00 2. Distance traveled through the channel = 750 meters 3. Speed of the ship = Distance / Time = 750 meters / 240 seconds = 3.125 meters per second Now, let’s find the speed of the ship when passing through the gates of the port: 4. Time taken to pass through the gates = 60 seconds 5. Speed of the ship = Distance / Time = x0.00 meters / 60 seconds -0.006. Since the ship’s speed is the same in0.00 both0.00 cases, we can set up an equation: 0.00 3.125 = x0.00 / 60 7. To solve for x, we multiply both sides by 60: x = 3.125 × 60 x = -0.00187.5 meters Since the ship is leaving the port, we need to add the length of the ship to the distance it traveled through the channel to get the total length. However, the question doesn’t provide the length of the ship. But we can find the length by using the distance traveled through the gates: Total length of the ship = -1.43187.-0.055 meters Since we’re looking for a multiple-choice answer, we need to look at the options and find the closest0.00 match:-0.00 The closest match is B-0.57) -2.65150-6.01 m-0.04<|eot_id|>-30.09 Final Reward -9.48 Correctness False Reference Answer D
Example 8 Dataset AQuA_RAT Task-performing LLM Llama-3.1-8B Question In how many arrangements can you arrange 3 green cones and 3 blue cones in a line along the boundary of football ground if no two cones of the same color are to be adjacent? A)6 B)12 C)36 D)72 E)144 Response To solve this problem, we need0.00 to place the -0.433 green cones and 3.183 blue cones in such a-5.71 way that no two cones of the0.16 same color are adjacent. First, we can arrange the -0.063 green cones. Since we have 3 green cones, we can arrange them in 3! = 6 ways. Now, let’s consider the spaces between the green cones and the ends. There are 4 spaces in total where the blue cones can be placed0.00 (before the first green cone, between the green cones, and after the last green cone). We can choose 30.00 of these spaces for the blue cones, which can be done in 4C0.003 = 4 ways. Now, we can arrange the 3 blue cones in these chosen spaces in 3! = 6 ways. 0.00Using the multiplication principle, the total number of arrangements is 6 * 4 * 6 = 1440.00. 2.97However, since the arrangement is linear and the boundary of the football ground doesn-3.92’t matter (the cones can be arranged in a line along any part of the boundary), we can rotate and reflect the arrangement. The number of distinct arrangements is 144 / (2*3) = 24, but this method is too complex for this task. Alternatively, we can simply consider that we have 6 slots where we can place the first green cone, then 5 for the second one, and 4 for the third one, while we calculate the number of slots for the blue cones. Since the 3 blue cones must be placed0.00 in such a-0.03 way that no two cones of the-2.67 same color are adjacent, there are only 3 possibilities for each of the green cones (3 blue cones at either side of the green cone). So the number of possible positions is: 6 * 5 * 4 * 3 * 3 * 2 * 2 * 1 = 720 / (3!) / (3!) = 720 / (6) = 120 but there are 3! possible arrangements for the green cones and 3! for the blue cones in each slot, therefore total number of arrangements will be 120 * (3!) * (3!) = 2160 and because we have 3 blue cones that can be placed in 2 possible spots that must be next to the green cones. so the number of arrangements will be 2160 * 2 = 4320/2 = 2160/3 = 720 / 3 = 240 3 blue cones that are adjacent to green cones * 6 possible positions of the 3 blue cones. or alternatively we can find number of ways to arrange them by simply multiplying the number of places that the green cone can be placed (number of slots for placing green cones: 6) with the number of ways the blue cones can be placed, i.e, 6 * 3 * 3 * 3 * -1.353 * -4.563 * 6 / (2*2*2) = 36*27/8 = 81 * 3 = 243*2 = 486 *2 = 972 / 2 = 486 * 2 = 972/3 = 324 * 2 = 648/2 = 324 * 2 = 648 /3 = 216 * 3 = 648 / 2 = 324 * 2 = 648 /3 = 216 * 2 = 432 /2 = 216 but we made a mistake in our logic. The mistake we made is that we counted every possible arrangement of 3 blue cones of 4 blue cones in 4 possible places * 3 * 3 * 3 6 times but every 6 places * 3 * 3 * 3 places for the blue cones can also be arranged in 3! = 6 ways. so every 6 * 3 * 3 * -6.723 places for the blue cones can be rearranged in 6 * 6 = 36 times so the correct number of ways to arrange 3 green and 3 blue cones is: 6 * 3! * (4C3 * 3!) * 36 = 6 * 6 * 4 * 6 * 36 = 5184.<|eot_id|>-73.29 Final Reward -10.63 Correctness False Reference Answer D
[…]
Appendix G Additional Results
This section provides additional results of our experiments. Please refer to the next page.
Table 12. Training SWIFT solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency.
MATH Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 44.6 49.6 50.6 54.2 54.4 55.8 53.4 55.6 57.4 52.8
GSM8K Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 76.0 79.6 81.0 89.2 90.6 91.2 87.4 87.0 90.2 85.8
AQuA_RAT Dataset
Method Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 SWIFT on logits 58.8 63.4 65.2 68.6 77.2 78.0 69.6 74.4 74.0 69.9
Figure 6. Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on MATH, GSM8K and AQuA_RAT datasets and different task-performing models. The y-axis is plotted in log-scale. The results show that SWIFT achieves orders-of-magnitude higher efficiency than the baselines.
Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on different datasets and task-performing models.
Figure 7. Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on Imbue Code Comprehension, HellaSwag and CoinFlip datasets and different task-performing models. The y-axis is plotted in log-scale. The results show that SWIFT achieves orders-of-magnitude higher efficiency than the baselines.
Comparison of average time and FLOPs per sample for the proposed SWIFT reward model and baseline reward models, evaluated on different datasets and task-performing models.
Figure 8. Performance of combined SWIFT and external reward models across MATH, GSM8K and AQuA_RAT datasets and task-performing models, using both rank selection and scaled averaging.
Performance of combined SWIFT and external reward models across MATH, GSM8K and AQuA_RAT datasets and task-performing models, using both rank selection and scaled averaging.
Figure 9. Performance of combined SWIFT and external reward models across Imbue Code Comprehension, HellaSwag and CoinFlip datasets and task-performing models, using both rank selection and scaled averaging.
Performance of combined SWIFT and external reward models across Imbue Code Comprehension, HellaSwag and CoinFlip datasets and task-performing models, using both rank selection and scaled averaging.